### v1

# Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

Abstract
Large language model-generated code (LLMgCode) has become increasingly prevalent in software development. Many studies report that LLMgCode has more quality and security issues than human-authored code (HaCode). It is common for LLMgCode to mix with HaCode in a code change, while the change is signed by only human developers, without being carefully checked. Many automated methods have been proposed to detect LLMgCode from HaCode, in which the perplexity-based method (Perplexity for short) is the state-of-the-art method. However, the efficacy evaluation of Perplexity has focused on the detection accuracy. In this article, we are interested in whether Perplexity is good enough in a wider range of realistic evaluation settings. To this end, we devise a large-scale dataset that includes 11,664 HaCode snippets and 13,164 LLMgCode snippets, and based on that, we carry out a family of experiments to compare Perplexity against feature-based and pre-training-based methods from three perspectives: (1) detection accuracy in terms of programming language, degree of difficulty, and scale of solution, (2) generalization capability, and (3) inference efficiency. The experimental results show that Perplexity has the best generalization capability while it has low accuracy and efficiency in most cases. Based on the experimental results and detection mechanism of Perplexity, we discuss implications into both the strengths and limitations of Perplexity, e.g., Perplexity is unsuitable for high-level programming languages while it has good interpretability. As the first large-scale investigation on detecting LLMgCode from HaCode, this article provides a wide range of evidence for future improvement.
Index Terms:
Code authorship attribution, human-authored code, LLM-generated code, large language model, perplexity
I Introduction
Large language model-generated code (LLMgCode) has become increasingly common in daily software development [1, 2]. Although LLMgCode reduce manual coding efforts and help developers focus on high-level tasks [3], the quality and security of LLMgCode are challenging [4, 5, 6, 7]. According to Pearce et al. [4], 40.73% of the 1,689 programs generated by GitHub Copilot are vulnerable, with risks far more than human-authored code (HaCode) [8, 9]. Moreover, LLMgCode has the risks of plagiarism and copyright infringement [10], which are less in HaCode. According to Guo et al. [11], LLMgCode can be a copy of HaCode even if the latter is with copyright protection [12]. Due to this, GitHub Copilot is being accused of compensating \$9 billion for generating restrictive code without attaching the necessary copyright or licenses [13].
In addition to the challenges of LLMgCode itself, the mixture of LLMgCode and HaCode has two challenges. The first is compromising the transparency of software development, which results in inaccurate attribution of code ownership in bug accountability and obscured workload assessment in creating software artifacts [14]. The second is increasing the difficulty of evaluating the reliability of third-party software, which results in distrust for upstream dependencies in the software supply chain [15]. As the authorship of LLMgCode has not been provided, detecting LLMgCode can be a stop-gap [15]. To this end, many studies [14, 16] and commercial tools [17] have proposed to detect LLMgCode from LLMgCode, which supports mitigating various risks by taking measures such as conducting stricter security inspections and performing code plagiarism detection [18].
The past few years have seen many studies [14, 15, 19, 20, 21, 22, 23, 24, 25] that proposed methods to detect LLMgCode from HaCode. According to the detection mechanisms used, these methods can be divided into three categories: 1. feature-based, 2. pre-training-based, and 3. perplexity-based methods (Perplexity for short for the last category). Feature-based methods [15, 19] are generally conventional machine learning classifiers that require professionals to devise appropriate features from aspects such as lexical and syntactic characteristics. Pre-training-based methods [20] take the code snippets as inputs. They convert code snippets to tokens, embed them as vectors, and use the pre-trained model as a classifier. Perplexity  [^1] For succinctness, Perplexity refers to the family of perplexity-based methods throughout the paper unless it is specifically stated otherwise. [14, 21, 22, 23, 24, 25] is a class of heuristic methods, which assume that LLMgCode has a lower perplexity than HaCode. Technically, perplexity is a measure of how confident a probability model predicts a sample [26]. When it comes to code generation, perplexity is used to measure the predictability of the next token of code based on prompts and existing tokens. The calculation of the perplexity is based on an LLM.
Figure 1: An example of calculating the perplexity of existing code
Figure 1 presents an example of calculating the perplexity of a probability model (CodeLlama) for code snippets (either LLMgCode or HaCode). The prompt used is shown in lines 2-3, and the code snippets to be calculated are in line 6. Based on the prompt, LLM calculates the probability of “  ” (Tab) as the first token is 0.20. Based on the prompt and “  ”, the LLM calculates the probability of “return” as the second token is 0.31. Followed by “ tf” and “.train” are 0.98 and 0.99, respectively. The higher the probability of a token, the more confident the model is about generating this token. Therefore, the perplexity of the model in understanding this token is low. Perplexity is a probability-based measure and has many implementations (e.g., log of probability [21], log of rank [23], and log of probability divided by log of rank [25]) in which the log of probability is common. For example, the perplexity of CodeLlama for “return” is $ln(0.31)=-1.18$. The perplexity of a model for a piece of code is the average perplexity of tokens.
When the code is natural and coherent, the model is confident in predicting the next token based on the prompts and the existing tokens, resulting in a lower perplexity. According to the mechanism of LLMs for code generation, LLMs generate code in the unit of a token through understanding previous tokens, calculating the probability of each possible next token, and applying the sampling strategy to select one with high probability as the output. Compared with LLMgCode, HaCode is more casual and personalized, making it more unpredictable. Therefore, it is reasonable that LLMgCode has lower perplexity than HaCode. Due to its explainability and scalability (e.g., log of probability, log of rank), Perplexity has been widely used [21, 22, 23, 24, 25].
Although Perplexity is becoming the most popular method, the efficacy evaluation of Perplexity is limited to detection accuracy. For example, Shi et al. [14] evaluated their proposed DetectCodeGPT using AUC (Area Under the Curve) on Python only. Yet it is unclear the accuracy (effectiveness) of DetectCodeGPT other than Python. DetectCodeGPT detects only the code generated by LLMs with fewer than 7B parameters, its effectiveness on the commonly used commercial models such as GPT-4o and Gemini is unclear. Moreover, DetectCodeGPT uses the same model for code generation and perplexity calculation. As it is hard to early access models for code generation, the effectiveness of DetectCodeGPT in a wide range of contexts is unclear. The limited evaluation of Perplexity does not reflect its effects well in practice.
Toward realistic evaluations of Perplexity, we conduct a large-scale investigation on Perplexity from three perspectives: detection accuracy, generalization capability, and inference efficiency. Detection accuracy is the foundation of detecting LLMgCode, while it fluctuates between situations, e.g., an effective method for C++ may not suit Python. Therefore, we evaluate the accuracy in various situations from three aspects. Specifically, the detection accuracy (effectiveness) is evaluated from the aspects of the programming language (C, C#, C++, Go, Java, Python, Ruby, and Rust), degree of difficulty (easy, middle, hard), and scale of solution (small, middle, and large). As new powerful LLMs are emerging, developers utilize various LLMs to assist in programming, therefore the generalization capability of Perplexity is important. The generalization capability is evaluated from the aspect of code generation models (GPT-3.5, GPT-4o, Gemini, and Llama). The waiting time greatly affects the user experience, thereby the inference efficiency of Perplexity should be considered. The inference efficiency is evaluated from the aspect of time spent (seconds/sample). We hope the multi-perspective efficacy evaluations can reflect the effects of Perplexity in practice. With experimental evaluation, we first devised a dataset containing 729 programming puzzles. Then, we collected 11,664 HaCode solutions from CodeNet [27] and generated 11,664 LLMgCode solutions using GPT-3.5. Next, we compared eight perplexity-based methods against two feature-based methods and one pre-training-based method that have been used to detect LLMgCode from HaCode.
The experimental results show that (1) for effectiveness, the best-performing Perplexity performs well for detecting LLMgCode in C/C++ (AUC=91%) and large-scale code (AUC=88%) but still does not perform as well as the other two methods regarding programming language, degree of difficulty, and scale of solution; (2) for generalization capability, Perplexity perform better than others across different LLMs, with AUC ranging from 80% to 82%; (3) for inference efficiency, Perplexity without perturbation is as efficient as others, while Perplexity with perturbation have a relatively slow efficiency, with detection speed ranging from 10 to 150 seconds per sample. Based on the experimental results, we provide some implications of Perplexity from both positive and negative perspectives. On the negative side, Perplexity is not suitable for high-level programming languages, short code snippets, and just-in-time detection. On the positive side, Perplexity has good interpretability, exhibits good generalization capability regarding LLMs, and has the potential for fine-grained detection.
The main contributions of this article can be summarized as follows.
  • Experimental Design: An experimental design to investigate the detection accuracy, generation capability, and inference efficiency of Perplexity for detecting LLMgCode.
  • Evidence: A wide range of evidence that illustrates advantages and limitations of Perplexity for detecting LLMgCode.
  • Dataset: A large-scale dataset for continuous research on detecting LLMgCode.
The remainder of this article is organized as follows. Section II presents background and related work. Section III illustrates motivating examples. Section IV and Section V elaborate on the research methodology and results, respectively. Section VI discusses implications. Finally, we present limitations in Section VII and conclude this article in Section VIII.
II Background and Related Work
This section gives an overview of LLMs for code generation and reviews studies that apply Perplexity for code authorship attribution.
II-A Background: Large Language Models for Code Generation
Since LLMs were proposed, they have been widely applied in many tasks across a variety of domains [28, 29, 30, 31]. One of the most exciting applications of LLMs is code generation [32, 33, 34, 35, 36]. LLMs are capable of generating source code based on the prompts. The prompts consist of natural and programmatic descriptions of code functions, retractions, input and output formats, or any other elements. Following the prompts, an LLM generates tokens sequentially until it arrives at a stop indicator or arrives at the token limit.
Many studies show that LLMgCode helps save time [37], reduce effort [38], and improve productivity [39] in software development. Moreover, LLMgCode help ease developers as they sometimes suffer from heavy workload [40]. With LLMgCode, developers can pay more attention to important and urgent tasks. Because of these benefits, LLMgCode has been widely used and even become a must-have in modern software development [41, 42]. A survey made with 500 USA developers showed that 92% of respondents use LLMgCode for both official assignments and personal use [43].
Although LLMs bring a lot of benefits, the quality and security of LLMgCode remain challenging [4, 44, 45]. One of the causes is the corpus used for training LLMs is not qualified and secure enough [5]. Therefore, LLMgCode might be buggy and vulnerable [46, 47]. For example, Fu et al. [48] assessed the security of LLMgCode in six common programming languages and found that 156 out of 435 code segments contained security vulnerabilities. The proportion of vulnerabilities in C++ and Go code accounts for 46.1% and 45.0%, respectively. Mastropaolo et al. [49] assessed the robustness of LLMgCode and found that slight modifications to the natural language description while still maintaining semantic equivalence would result in the LLMs generating different results, occurring in 46% (408/892) of the total samples. Jimenez et al. [50] assessed the capability of large language models to solve 2294 real-world problems (Issues) on GitHub and found that the pass rate of code generated using GPT-4 has the lowest pass rate–0%, using Claude 2 has the highest pass rate–1.96%. In addition, LLMgCode might be similar or even unchanged to existing code snippets in the training corpus [11]. As the training corpus might include a significant amount of restrictive code [12], there are copyright infringement risks for LLMgCode.
II-B Related Work: Perplexity for Code Authorship Attribution
Detecting LLMgCode from HaCode can be regarded as a particular code authorship attribution problem [51, 52, 53, 54] in which the authors are either LLMs or humans. Code authorship attribution fills in the missing or inaccurate author information of source code fragments, which is beneficial for software maintenance [55, 56, 57] and software quality analysis [58, 59, 60, 61]. For example, code authorship attribution helps detect code plagiarism [62, 63] and trace malware provenance [64]. Moreover, detecting LLMgCode from HaCode increases the authenticity and integrity of software [14], reducing risks such as security vulnerabilities and copyright infringement in the software supply chain [15].
Perplexity is originally one of the most popular methods for detecting AI-generated content (including but not limited to LLMgCode) [23, 24, 25]. For example, Mitchell et al. [24] proposed DetectGPT to detect LLM-generated text. DetectGPT utilizes the log probabilities to calculate the perplexity. The authors evaluated DetectGPT using the AUC metric on 500 LLM-generated samples and 500 human-written samples. In addition, they evaluated the generalization capability of DetectGPT by alternately using GPT-J-6B, GPT-Neo-2.7B, and GPT-2-1.5B as the code generation model and the perplexity calculation model. Su et al. [25] proposed two methods to calculate the perplexity. The first is Log-Likelihood Log-Rank Ratio (LRR) and the second is Normalized Perturbed Log-Rank (NPR). The authors evaluated LRR and NPR using the AUC metric on 300 LLM-generated and 300 human-written samples. They also compared LRR and NPR against baselines in terms of inference efficiency.
With the increasing popularity of LLMgCode, Perplexity has been commonly used for detecting LLMgCode from HaCode [14, 26]. For example, Xu and Sheng [26] proposed a method using the variant of probability to calculate the perplexity. They evaluated their methods using the AUC, FPR, and TPR metrics on 5,214 LLM-generated and 5,214 human-authored code snippets. They also evaluated the influence of models used to calculate perplexity. Shi et al. [14] proposed a method namely DetectCodeGPT to detect LLMgCode from HaCode. DetectCodeGPT calculates the perplexity based on NPR. They evaluated DetectCodeGPT using the AUC metric on 500 LLM-generated and 500 human-authored code snippets. They also evaluated the generalization capability of DetectCodeGPT by alternately using Incoder-1B, Phi-1B, StarCoder-3B, WiardCoder-3B, CodeGen2-3.7B, and GodeLlama-7B as the code generation model and the perplexity calculation model.
Overall, Perplexity is a popular method for detecting LLMgCode from HaCode but its efficacy evaluation is limited, which does not reflect its effects well in practice. In this article, we carry out a large-scale investigation on Perplexity from multiple realistic aspects.
III Motivating Examples
This section describes three examples to motivate this investigation.
III-A Code Authenticity
Although Perplexity has been demonstrated useful for detecting LLMgCode from HaCode, most of the LLMgCode is generated by LLMs with relatively a few parameters, such as CodeLlama-7B [14].
Figure 2: An example of the code generated by CodeLlama-7B and the code written by human
Figure 2 presents examples of LLMgCode and HaCode. The prompt indicates that the function to be generated is to make a git commit when the template files are modified. The prompt is the input of CodeLlama-7B (shown in the first block). The LLMgCode is the solution generated by CodeLlama-7B (shown in the second block), and the HaCode is the solution provided by a human developer (shown in the third block). The LLMgCode is not an acceptable solution and differs significantly from the HaCode. Specifically, the LLMgCode contains only repeated comments. Although it is easy to detect such LLMgCode from HaCode, few developers use such a code generation model like CodeLlama-7B as it is far from addressing real-world problems.
When Perplexity uses the same model for code generation and detection, it might achieve low perplexities of LLMgCode. This is because the model would assign a high prediction probability to each token when it perceives the code generated by itself. However, in real-world scenarios, the detection model does not have clues to know which model is used to generate code early. Moreover, these LLMs used for generating code have hundreds of billions of parameters, making them challenging to deploy locally to calculate perplexities.
III-B Language Diversity
Perplexity is evaluated in limited programming languages so far. However, the code stylometry of different programming languages varies greatly. Figure 3 presents two pieces of code generated by GPT-3.5, both of which are to calculate the median of all subarrays of a given array, and then calculate the median of these medians.
Figure 3: An example of code achieving the same task in Python and C
As illustrated in Figure 3, the code stylometries differ in Python and C, particularly in their logic and scale. Python has a rich standard library and built-in functions. For example, “statistics.comedian” is a tool library in Python. It can directly calculate the median without the need for manual sorting and selection. while C does not have a direct median calculation function, so we need to use “qsort” to sort and manually calculate the median. In fact, many high-level functions must be manually implemented in C, such as inserting or deleting elements in arrays. However, such functions can be automatically handled by the built-in methods of lists in Python. In addition, Python is a dynamically typed programming language that does not require declaring variable types, making the code more concise. While C is a statically typed programming language where each variable needs to declare its type. Moreover, manipulating elements such as list slicing is easy in Python. For example, “a[i: j+1]” returns subarrays without manual calculation and traversal. While C requires manually managing memory and arrays, which means that developers need to handle each element carefully.
Due to differences such as the built-in library support and design principles, code written in different programming languages exhibits variations. In addition to programming languages, the code used to solve puzzles of different difficulty and the code of different scales exhibits variations in terms of logic and complexity.
III-C Generalization Capability
There are many LLMs available for generating code, such as GPT-4o, Copilot, and Gemini. Due to differences in training data, model architecture, optimization objectives, generation strategies, etc., the code generated by these LLMs is different. Figure 4 presents two Java code snippets generated by GPT-4o and Gemini solving the same programming puzzle. The puzzle is a dynamic programming problem that is similar to the unbounded knapsack problem. This problem can be described as follows. Given the monster’s initial health $H$ and $N$ different spells, each spell has a fixed damage $A_{i}$ and a magic point cost $B_{i}$. The objective is to reduce the monster’s health to 0 or below by casting spells while minimizing the total magic points consumed.
Figure 4: An example of code achieving the same task generated by GPT-4o and Gemini
As shown in Figure 4, although both pieces of code aim to reduce the monster’s health to 0 with minimal loss, the two LLMs have different implementations. GPT-4o starts to generate code from the perspective of damage, aiming to achieve total damage $H$ with the minimum cost. Gemini starts to generate code from the perspective of the monster’s health, aiming to reduce the monster’s health to 0 with the minimum cost. The former is to intuitively accumulate damage in the forward direction, and the latter is to reduce the monster’s health in the reverse direction.
In addition to the above differences, the abstraction levels of implementations are different. The solution generated by Gemini is detailed and complex. Gemini uses a two-dimensional array ‘dp’ to simultaneously record health values and spell combinations. GPT-4o only uses a one-bit array ‘dp’ to record the minimum cost of achieving each damage value. When updating the state, Gemini considers all spell combinations in each update, making it more complex; while GPT-4o directly updates the state by traversing damage values and spell methods. In addition, Gemini considers boundary conditions and special cases, such as returning an outlier of “-1” when the target cannot be reached. Unfortunately, GPT-4o fails to handle these boundary conditions, reducing its robustness.
In short, GPT-4o focuses on the core logic; while Gemini focuses on practical usages, boundary conditions, and robustness on this example. This example motivates us to investigate the generalization capability of Perplexity when the code is generated by different LLMs. In addition to generalization, the efficiency of Perplexity is important, as it influences the usage scenarios and user experience of Perplexity.
IV Research Methodology
This section elaborates on the research methodology of detecting LLMgCode from HaCode. shows an overview of our research design and process.
IV-A Research Questions
Existing work focuses on evaluating the detection accuracy (effectiveness) of Perplexity, yet there are many aspects should be considered when using it in practice. Toward realistic detection, we intend to answer three research questions (RQs) through this investigation.
  • $\text{{RQ}}_{\textbf{1}}$:
    Is Perplexity an effective method for detecting LLMgCode?: Language ($\text{{RQ}}_{\textbf{1.1}}$), Difficulty ($\text{{RQ}}_{\textbf{1.2}}$), and Scale ($\text{{RQ}}_{\textbf{1.3}}$)
    Motivation:
    Detection accuracy is crucial for detecting LLMgCode, while it might not be stable between situations, e.g., an effective method for C++ may not suit Python. Therefore, we propose $\text{RQ}_{1}$ to compare Perplexity against state-of-the-art methods regarding effectiveness from three sub-aspects: (1) programming languageof the implementation: C, C#, C++, Go, Java, Python, Ruby, and Rust; (2) degree of difficultyof the puzzle: easy, medium, and hard; and (3) scale of solutionto the puzzle: small, middle, and large.
  • $\text{{RQ}}_{\textbf{2}}$:
    Does Perplexity have the generalization capability?
    Motivation:
    Developers utilize various LLMs to assist in programming. Due to differences in training data, model architecture, and generation strategies, the code generated by these LLMs can vary significantly. An optimal detection method should be capable of detecting the code generated by different LLMs. In $\text{RQ}_{2}$, we investigate the generalization capability of Perplexity in detecting LLMgCode.
  • $\text{{RQ}}_{\textbf{3}}$:
    Is Perplexity an efficient method for detecting LLMgCode?
    Motivation:
    Detection speed (efficiency) influences the user experience. According to the detection mechanism, Perplexity applies perturbation strategies to generate dozens of perturbation samples and uses an LLM to calculate the probability or rank of each token for both the original sample and perturbation samples [14, 25]. Such a two-phase detection process might be time-consuming. In $\text{RQ}_{3}$, we compare Perplexity with baselines from the perspective of inference efficiency.
IV-B Data Preparation
Although there are several dataset [14, 15, 19] for detecting LLMgCode from HaCode, none of them is suitable to address all research questions. Therefore, we developed a new one based on CodeNet [27]. CodeNet was originally proposed by IBM and built to promote AI technologies in the domain of code. Since then, CodeNet has been used for many studies [26, 65, 66, 67, 68, 69].
Figure 5: An example of a programming puzzle in CodeNet
As shown in Figure 5, a programming puzzle provided by CodeNet has seven basic elements: (1) statement, (2) input, (3) output, (4) constraint, (5) difficulty score, (6) sample input, and (7) sample output. The statement describes what the puzzle is. The input and output describe the requirements for input and output, respectively. The constraint describes the specific constraints on the variables. The difficulty score measures the difficulty level of the puzzle. Finally, the sample input and sample output provide the input and output of a specific example, respectively. Based on CodeNet, we developed our dataset through the following five steps.
IV-B1 Step-1: Problem Selection
To build a standard dataset for evaluation, we excluded the programming problem whose (1) statement is not in English as it is the most popular language; (2) statement contains images or tables as they can hardly be extracted and input into LLMs; (3) solutions are all less than 30 LOC (Lines Of Code) as they are mainly for trivial puzzles; (4) solutions are incomplete regarding all the eight programming languages as they constitute the HaCode to be evaluated in $\text{RQ}_{1.1}$; (5) structure is not complete, i.e., statement, input, output, constraints, difficulty, and samples ($\geq 2$). After employing these criteria, our dataset includes 729 programming puzzles.
IV-B2 Step-2: HaCode Collection
We randomly collected two solutions that were provided by human developers from CodeNet for each puzzle regarding each of the eight programming languages.
IV-B3 Step-3: LLMgCode Generation
We generated LLMgCode for programming puzzles using OpenAI GPT-3.5 (gpt-3.5-turbo-0125). The prompts for code generation are shown in Figure 6. There are two solutions for each puzzle regarding each of the eight programming languages. Moreover, to address $\text{RQ}_{2}$, we used the prompts from Figure 6 and input them into GPT-4o, Gemini, and Llama, with each model generating 500 solutions. Note that the following two steps are performed on the code solutions that have been generated by GPT-3.5.
Figure 6: The prompts used for generating LLMgCode solutions for programming puzzles
IV-B4 Step-4: Classification According to the Difficulty of Puzzles
We classified the puzzles and solutions into three groups based on difficulty scores of puzzles: easy for $\leq 200$, medium for $>200$ and $\leq 500$, and hard for $>500$.
IV-B5 Step-5: Classification According to the Scale of Solutions
We classified the puzzles and solutions into three groups based on the scale of solutions: small for $\leq 20$, medium for $>20$ and $\leq 50$, and large for $>50$.
IV-C Studied Methods
We compared Perplexity against feature-based and pre-training-based methods aimed at detecting LLMgCode from HaCode as the experimental subjects.
IV-C1 Feature-based Methods
Two feature-based methods are as follows.
  • Idialu et al. [19] that leverages a combination of layout, syntactic, and lexical features in Python source code as the input. This method uses XGBoost [70] to construct the detection method.
  • Bukhari et al. [15] that takes lexical, syntactic features, and their combination in C source code as the input. This method uses Random Forest, SVM, KNN, and XGBoost to build the detection methods.
IV-C2 Pre-Training-based Method
The only one pre-training-based method is as follows.
  • GPTSniffer [20] that transforms the source code into a sequence of independent tokens. These tokens were processed by the pre-trained model, i.e., CodeBERT. This method uses CodeBERT to develop the detection method.
IV-C3 Perplexity
Eight popular perplexity-based methods are as follows. The first five calculate the perplexity based on the original sample without perturbations. The last three require applying other LLMs to perturb the original sample and calculate the perplexity based on the discrepancies before and after the perturbation.
  • Log-p(x) [21] that utilizes the average token-wise log probability to calculate the perplexity. The log probability is the logarithm of the probability of the next token (ground truth) based on previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a larger log-p(x) than HaCode.
  • Entropy [22] that uses the average entropy to calculate the perplexity. Entropy is the sum of the probability multiplied by the negative logarithm of the probability of each token. This method assumes that LLMgCode has a larger entropy than HaCode.
  • Log-Rank/Rank [22, 23] that utilizes the average log-rank/rank of each token to calculate the perplexity. Log-Rank/Rank is the logarithm of rank/rank of the next token in the vocabulary list. The vocabulary list is sorted according to the probability of each token given the previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a lower log-rank/rank than HaCode.
  • LRR [25] that divides log-likelihood by log-rank to calculate the perplexity. Log-likelihood is the average token-wise log probability, representing the absolute confidence for the true token. Log-rank is the average token-wise log rank, representing the relative confidence for the true token. For LLMgCode, log-rank is more discernible than the log-likelihood. This method assumes that LLMgCode has a larger LRR than HaCode.
  • DetectGPT [24] that leverages the discrepancy between the log probability of the original sample (a collection of tokens) and the average log probability of the perturbed samples to calculate the perplexity. The perturbed samples are obtained by masking and fulfilling some tokens. This method assumes that LLMgCode has a greater perturbation discrepancy as it is more sensitive to perturbations than HaCode.
  • NPR [25] that divides the average log-rank of the perturbed samples by the log-rank of the original sample to calculate the perplexity. The perturbed samples are obtained by replacing some tokens. This method assumes that LLMgCode has a larger NPR as it is more susceptible to perturbations than HaCode.
  • DetectCodeGPT [14] that utilizes the discrepancy between the NPR of the original sample and the average NPR of the perturbed samples to calculate the perplexity. The perturbed samples are obtained by adding spaces and new lines to the original sample. This method assumes that the average NPR of perturbed LLMgCode drops more than HaCode.
IV-D Experimental Settings
Following related work [14, 19, 24, 25], we use the AUC (Area Under the Curve) to measure each detection method. AUC stands for the area under ROC (Receiver Operating Characteristic curve). Given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the calculation of AUC is as follows.
[…]
where $t$ denotes the varying threshold values.
For $\text{RQ}_{1.1}$, we evaluated Perplexity and pre-training-based methods in all the eight programming languages. We evaluated feature-based methods on specific languages because those features are carefully designed for those languages. Due to the time and cost limit, we tested 500 pairs of human-LLM code snippets for each aspect regarding each research question. For $\text{RQ}_{1}$, we tested 500 pairs of human-LLM code snippets for each programming language, difficulty level, and scale. For $\text{RQ}_{2}$ and $\text{RQ}_{3}$, we tested 500 pairs of human-LLM code snippets for each LLM. We randomly selected 500 pairs of human-LLM code snippets from the dataset each time. Following Shi et al. [14], we utilized CodeT5-770M to generate perturbed code snippets for DetectGPT. We added spaces and empty lines to generate perturbed code snippets for NPR and DetectCodeGPT. We generated 50 perturbed samples for each code sample. We carried out the experiments using 1 NVIDIA A100-PCIE-40GB.
V Results and Analysis
This section reports on results and analysis to answer research questions.
V-A $RQ_{\text{1.1}}$: Programming Language
Table I and Table II shows the AUC of Perplexity, and feature-based and pre-training-based methods from the perspective of programming languages, respectively. As the feature-based detection methods are designed for specific languages, we provide only the results in Python (Idialu) and C (Bukhari) in Table II. For presentation clarity, we highlight notable results with blue backgrounds (AUC:$90\%-100\%$;$80\%-90\%$;$70\%-80\%$).
TABLE I: The AUC of Perplexity in different programming languages
Lan Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
C Incoder 56.43 38.55 19.07 59.98 64.80 57.31 84.98 84.80
Phi-1 83.66 36.19 59.41 77.13 25.84 61.66 90.17 90.40
StarCoder 65.43 48.50 49.85 66.72 58.72 56.32 86.47 85.70
CodeGen2 74.30 39.72 52.23 72.12 53.33 56.40 88.51 87.08
WizCoder 80.87 30.02 58.61 77.82 53.16 58.65 88.73 89.09
CodeLlama 74.75 41.99 57.49 70.25 58.83 57.27 85.05 84.95
Average 72.57 39.16 49.44 70.67 52.45 57.94 87.32 87.00
C# Incoder 48.75 48.97 53.59 56.97 82.05 33.17 64.66 65.12
Phi-1 90.12 41.50 84.61 86.90 36.59 39.28 84.54 86.11
StarCoder 62.09 55.50 72.55 67.69 71.59 27.34 73.39 73.62
CodeGen2 75.45 45.08 77.50 77.80 68.86 35.67 78.85 78.55
WizCoder 76.73 36.49 75.41 76.48 54.82 33.22 76.45 79.04
CodeLlama 62.96 59.45 72.53 67.66 71.32 35.06 66.94 68.34
Average 69.35 47.83 72.70 72.25 64.21 33.96 74.14 75.13
C++ Incoder 63.99 34.96 19.79 67.89 69.92 56.17 81.90 82.38
Phi-1 88.38 35.20 63.85 82.41 34.02 59.22 91.14 92.10
StarCoder 75.17 44.86 56.75 74.79 70.28 54.83 89.56 89.35
CodeGen2 79.91 39.33 59.04 74.43 57.40 56.00 86.33 86.04
WizCoder 83.06 31.19 61.39 83.11 56.86 57.46 90.52 91.37
CodeLlama 76.00 48.37 59.64 77.39 67.68 56.51 86.08 86.65
Average 77.75 38.99 53.41 76.67 59.36 56.70 87.59 87.98
Go Incoder 41.91 53.58 22.65 43.97 71.24 29.42 54.83 54.47
Phi-1 72.68 45.99 61.24 68.27 40.24 37.04 68.74 70.10
StarCoder 50.95 60.61 52.13 49.62 47.88 32.19 53.33 53.59
CodeGen2 53.17 62.07 54.42 53.88 45.76 35.64 55.68 55.54
WizCoder 64.71 60.73 48.72 57.91 45.15 32.09 62.42 64.21
CodeLlama 47.84 65.36 53.63 49.98 49.25 32.54 53.00 54.40
Average 55.21 58.06 48.80 53.94 49.92 33.15 58.00 58.72
Java Incoder 54.61 44.73 21.18 61.72 67.74 42.03 75.11 74.65
Phi-1 85.44 40.40 62.55 79.07 28.80 49.80 84.69 85.54
StarCoder 63.01 56.90 55.61 61.27 53.33 42.25 74.87 73.67
CodeGen2 71.96 50.69 51.89 66.52 49.20 44.81 72.75 71.94
WizCoder 69.97 45.14 59.85 65.34 47.12 43.20 69.29 70.36
CodeLlama 64.27 57.67 54.88 61.21 56.86 41.67 65.65 65.95
Average 68.21 49.26 50.99 65.85 50.51 43.96 73.73 75.23
Python Incoder 47.33 51.23 44.73 50.34 57.62 41.02 60.91 58.52
Phi-1 60.03 53.48 48.46 56.93 34.82 45.08 63.89 62.64
StarCoder 55.72 53.83 40.02 52.48 37.86 40.96 58.03 55.70
CodeGen2 56.54 53.36 42.20 50.80 34.18 38.70 59.02 56.48
WizCoder 62.79 44.41 47.66 56.25 23.82 42.54 63.61 62.52
CodeLlama 57.11 48.76 43.67 48.84 32.16 40.25 55.94 54.60
Average 56.59 50.84 44.46 52.61 36.74 41.43 60.23 58.41
Ruby Incoder 52.41 44.31 46.83 51.24 61.11 39.17 55.49 54.16
Phi-1 67.47 39.48 63.86 65.58 54.34 44.45 66.32 65.64
StarCoder 61.61 44.25 47.51 59.77 42.90 38.70 60.04 57.52
CodeGen2 63.78 40.34 53.65 59.29 45.46 41.44 63.95 61.61
WizCoder 70.22 33.63 50.58 66.85 41.83 44.25 73.44 71.65
CodeLlama 61.63 42.48 57.81 65.62 58.55 39.15 73.17 72.05
Average 62.85 40.75 53.37 61.39 50.70 41.19 65.40 63.77
Rust Incoder 68.41 38.70 69.87 71.97 23.03 79.82 71.55 72.04
Phi-1 89.72 20.69 80.27 88.28 26.17 61.87 82.22 83.63
StarCoder 39.10 73.59 49.26 42.53 23.63 55.23 43.45 43.14
CodeGen2 66.15 54.97 58.30 66.04 23.94 55.24 63.06 63.15
WizCoder 64.41 51.38 57.72 57.91 27.22 36.61 55.29 55.96
CodeLlama 57.79 66.74 59.09 58.60 24.65 53.55 55.54 56.44
Average 64.26 51.01 62.42 64.22 24.77 57.05 61.85 62.39
  • 1
    For succinctness, we abbreviate the baseline names as follows: log-p (Log-p(x)), Ent (Entropy), L-Rank (Log-Rank), DGPT (DetectGPT), DCGPT (DetectCodeGPT).
TABLE II: The AUC of feature-based and pre-training-based methods in different programming languages
Lan Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
C / 99.01 99.13
C# / / 98.15
C++ / / 99.58
Go / / 94.19
Java / / 0.9611
Python 89.60 / 95.17
Ruby / / 93.25
Rust / / 98.64
When comparing Perplexity against the other two classes of methods, we select the best-performing method as representative [^2] For succinctness, we adopt this measure throughout the paper. Perplexity does not perform as well as GPTSniffer which is a pre-training-based method in all the eight programming languages. The AUC of Perplexity and GPTSniffer for each language are as follows: C (90.40%–99.13%), C# (90.12%–98.15%), C++ (91.37%–99.58%), Go (72.68%–94.19%), Java (85.54%–96.11%), Python (63.89%–95.17%), Ruby (73.44%–93.25%), and Rust (89.72%–98.64%). The difference can be explained that GPTSniffer is a supervised method while Perplexity is not. Moreover, Perplexity cannot compare to the feature-based methods in Python and C: Python (63.89%–89.60%), C (90.40%–99.01%).
When comparing the AUC of Perplexity between programming languages, we observe that Perplexity performs best in C/C++ in which NPR and DetectCodeGPT achieve an average AUC of 85%. However, NPR, DetectCodeGPT, and other methods cannot achieve an AUC of over 80% in other languages. Taking C/C++ and Python as examples. C/C++ requires developers to implement the majority of functions manually, which however are provided by third-party libraries in Python. In this case, C/C++ has more code stylometry than Python [64]. However, the design philosophy of Python encourages simplified code stylometry, making it easy for LLMs to generate human-style code.
$\textbf{RQ}_{\textbf{1.1}}$: Perplexity can not perform as well as the feature-based and pre-training-based methods from the perspective of programming languages. Among eight languages, Perplexity is only effective in C and C++.
V-B $RQ_{\text{1.2}}$: Degree of Difficulty
Table III and Table IV present the AUC of Perplexity, and feature-based and pre-training-based methods from the perspective of the degree of difficulty, respectively.
TABLE III: The AUC of Perplexity in different degrees of difficulty
Diff Code LLM without perturbation with perturbation
log-p Ent Rank LRank LRR DGPT NPR DCGPT
Easy Incoder 51.75 46.33 40.75 57.76 61.56 46.80 65.03 64.46
Phi-1 69.63 44.47 57.08 69.34 38.90 57.84 75.41 75.09
StarCoder 56.23 52.40 49.72 56.86 50.90 46.27 64.13 62.79
CodeGen2 62.19 46.78 53.82 59.32 49.77 45.71 65.50 64.02
WizCoder 67.84 41.60 55.04 59.52 39.82 47.72 64.72 64.31
CodeLlama 59.84 49.69 53.81 58.15 52.16 45.62 63.60 62.91
Average 61.25 46.88 51.70 60.16 48.85 48.33 66.40 65.60
Middle Incoder 53.30 45.55 41.57 57.04 66.77 42.51 66.02 65.56
Phi-1 76.54 39.33 65.22 71.44 42.02 46.18 76.29 77.00
StarCoder 55.35 56.54 52.41 54.32 49.97 43.12 61.36 60.56
CodeGen2 62.77 49.94 53.28 63.96 53.03 46.77 69.02 68.22
WizCoder 66.38 43.22 56.73 62.90 43.73 43.51 68.18 68.69
CodeLlama 57.67 54.05 52.91 60.93 55.40 41.20 66.08 65.97
Average 62.00 48.11 53.69 61.76 51.82 43.88 67.82 67.67
Hard Incoder 54.53 44.44 37.73 64.05 76.48 41.64 70.27 70.92
Phi-1 75.92 40.94 70.32 72.25 46.92 41.63 74.69 76.07
StarCoder 56.30 57.90 58.33 54.99 53.88 34.26 62.48 62.47
CodeGen2 65.25 50.16 61.54 67.26 57.51 39.70 71.23 71.21
WizCoder 66.28 41.05 63.58 66.32 49.05 36.55 69.21 70.53
CodeLlama 62.97 54.87 65.99 62.32 60.48 35.16 65.82 66.67
Average 63.54 48.23 59.58 64.53 57.39 38.16 68.95 69.64
TABLE IV: The AUC of feature-based and pre-training-based methods in different degrees of difficulty
Diff Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
Easy 85.37 96.40 90.57
Middle 89.60 98.98 94.32
Hard 93.18 98.25 94.33
Perplexity exhibits lower performance compared with the feature-based and pre-training-based methods in terms of all three levels of difficulty. The differences between Perplexity and pre-training-based methods for each level of difficulty are as follows: easy (75.41%–90.57%), middle (77.00%–94.32%), and hard (76.48%–94.33%). When compared with feature-based methods, Perplexity is less effective than Idialu: easy (75.41%–85.37%), middle (77.00%–89.60%), and hard (76.48%–93.18%), and Bukhari: easy (75.41%–96.40%), middle (77.00%–98.98%), and hard (76.48%–98.25%). The discrepancy can be explained that feature-based and pre-training-based methods are based on supervised training while Perplexity not.
When it comes to the AUC of Perplexity between different degrees of difficulty, we find that Perplexity performs not well in all three levels of difficulty, but slightly improves from the easy to hard level. To be specific, log p(x) and Log Rank are the best-performing methods in Perplexity without perturbation, achieving the AUC of 61.25% and 60.10% at the easy level, 62.00%, and 61.76% at the middle level, and 63.54% and 64.53% at the hard level, respectively. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve the AUC of 66.40% and 65.60% at the easy level, 67.82% and 67.67% at the middle level, and 68.95% and 69.64% at the hard level, respectively. Since those solutions to difficult puzzles have complex logic and customized implementations, they have more code stylometry than solutions to easy puzzles.
$\textbf{RQ}_{\textbf{1.2}}$: Perplexity perform not well in all three levels of difficulty, although the performance gradually and slightly improves from the easy to the hard level. The feature-based and pre-training-based methods perform well at all three levels of difficulty.
V-C $RQ_{\text{1.3}}$: Scale of Solution
Table V and Table VI show the AUC of three classes of detection methods, respectively. In Table VI, as there is no large-scale code solution in Python, we provide only the results in the small and middle scale for Idialu.
TABLE V: The AUC of Perplexity in different scales of solution
Scale Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
Small Incoder 52.94 45.41 47.20 52.79 53.91 47.55 58.46 57.24
Phi-1 65.07 42.43 53.11 62.33 42.86 60.14 69.75 68.63
StarCoder 61.72 43.41 49.76 53.42 41.05 54.93 59.29 56.89
CodeGen2 60.65 43.23 50.69 55.45 44.03 47.30 63.75 61.25
WizCoder 66.42 38.99 53.71 63.48 40.26 52.51 69.47 68.01
CodeLlama 61.68 40.38 54.39 60.45 50.67 43.50 67.95 66.35
Average 61.41 42.31 51.48 57.99 45.46 50.99 64.78 63.06
Middle Incoder 57.47 43.11 39.15 61.29 63.56 57.28 72.93 72.04
Phi-1 75.35 38.13 55.44 67.33 35.52 59.11 77.42 77.26
StarCoder 63.84 47.49 54.01 59.42 51.61 56.54 70.22 69.36
CodeGen2 68.56 41.39 51.73 63.92 44.98 56.55 72.55 71.29
WizCoder 72.77 34.01 52.63 63.36 38.64 57.51 73.97 73.64
CodeLlama 63.85 46.76 51.77 62.86 50.66 54.37 70.38 69.56
Average 66.97 41.82 50.79 63.03 47.50 56.89 72.91 72.19
Large Incoder 72.17 29.05 47.88 75.28 77.86 48.97 77.12 78.39
Phi-1 87.81 21.03 73.76 86.00 58.52 50.81 82.22 83.94
StarCoder 67.85 46.41 61.18 67.59 58.67 45.13 68.30 69.46
CodeGen2 78.21 35.07 65.32 77.19 61.70 43.76 75.46 76.62
WizCoder 82.75 21.61 71.10 81.27 57.75 47.70 81.04 82.24
CodeLlama 76.40 42.48 69.43 76.34 64.85 45.54 76.70 77.94
Average 77.53 32.61 64.78 77.28 63.22 46.99 76.81 78.10
TABLE VI: The AUC of feature-based and pre-training-based methods in different scales of solution
Scale Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
Small 85.87 91.20 82.74
Middle 92.58 98.26 96.10
Large / 96.67 94.19
Perplexity does not perform as well as the pre-training-based method, i.e., GPTSniffer in all the three scales. The AUC of Perplexity and GPTSniffer for each scale are: small (69.75%–82.74%), middle (77.42%–96.10%), and large (87.81%–94.19%). For the featured-based method, Perplexity is less effective than Idialu: small (69.75%–85.87%), middle (77.42%–92.58%), and Bukhari: small (69.75%–91.20%), middle (77.42%–98.26%), and large (87.81%–96.67%). These differences are mainly because feature-based and pre-training-based methods are supervised methods while Perplexity is a collection of heuristic methods.
When comparing the methods in Perplexity regarding different scales, we observe that those well-performed methods improve from the small to large scale. For example, log p(x) is the best-performing method in Perplexity without perturbation, with an AUC of 61.41% for small, 66.97% for middle, and 77.53% for large. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve an AUC of 64.87% and 63.06% for small, 72.91% and 72.19% for middle, and 76.81% and 78.10% for large, respectively. These results can be explained that large-scale code snippets have more diverse and detailed implementations than others, which makes HaCode more distinct from LLMgCode.
$\textbf{RQ}_{\textbf{1.3}}$: Perplexity performs less effectively than the feature-based and pre-training-based methods in all the three scales. Perplexity only performs adequately in the code with large scale.
V-D $RQ_{\text{2}}$: Generalization Capability
Table VII and Table VIII present the AUC of Perplexity and the other two classes of methods in detecting code generated by four LLMs.
TABLE VII: The AUC of Perplexity in different LLMs
Model Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
GPT-3.5 Incoder 51.89 46.86 39.93 63.01 72.28 49.68 73.10 72.78
Phi-1 75.44 41.93 66.19 75.87 45.27 52.79 81.10 81.64
StarCoder 56.14 54.82 55.86 61.09 57.55 47.53 70.86 70.03
CodeGen2 63.27 49.26 58.56 66.84 56.52 48.27 73.31 72.40
WizCoder 66.34 42.77 60.98 68.54 50.25 72.24 74.26 74.37
CodeLlama 59.41 53.78 60.53 65.79 60.39 41.48 71.52 71.28
Average 62.08 48.24 57.01 66.86 57.04 52.00 74.03 73.75
GPT-4o Incoder 55.29 44.88 44.15 63.16 63.12 54.10 74.89 74.71
Phi-1 72.89 38.02 63.79 72.56 44.10 58.14 80.52 80.88
StarCoder 56.04 52.69 53.07 58.90 53.02 51.63 71.25 70.44
CodeGen2 62.52 48.15 54.12 63.91 51.03 53.03 72.10 71.55
WizCoder 65.54 40.62 57.90 65.91 45.73 66.75 72.95 73.00
CodeLlama 60.70 50.93 56.88 64.68 54.86 48.01 71.13 70.93
Average 62.16 45.88 54.98 64.85 51.98 55.28 73.81 73.59
Gemini Incoder 54.53 45.43 42.55 63.70 76.00 51.37 73.74 73.17
Phi-1 74.55 42.25 71.10 74.30 51.08 54.88 81.78 81.66
StarCoder 60.58 53.99 64.18 64.75 66.84 49.24 73.36 72.26
CodeGen2 67.39 48.34 65.84 69.68 65.25 50.68 76.23 74.98
WizCoder 69.21 44.70 64.43 67.99 50.52 58.07 74.36 74.15
CodeLlama 63.76 53.00 68.01 66.36 68.45 43.83 73.49 73.21
Average 65.00 47.95 62.68 67.80 63.02 51.35 75.49 74.90
Llama Incoder 57.00 42.88 43.15 65.10 74.82 50.84 73.20 73.14
Phi-1 74.45 41.33 68.64 73.47 46.39 54.96 79.42 79.55
StarCoder 66.20 47.63 67.67 70.29 72.05 51.47 75.56 75.23
CodeGen2 71.42 42.79 66.51 73.05 66.92 49.79 77.42 76.95
WizCoder 74.69 38.59 68.99 76.39 56.50 67.27 79.08 78.74
CodeLlama 69.59 47.23 71.41 73.78 72.37 52.15 77.52 77.69
Average 68.89 43.41 64.39 72.01 64.84 54.41 77.03 76.88
TABLE VIII: The AUC of feature-based and pre-training-based methods in different LLMs
LLM Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
GPT-3.5 91.43 98.63 92.21
GPT-4o 83.16 98.49 90.30
Gemini 62.77 92.13 75.21
Llama 71.76 97.15 83.26
Perplexity has a better generalization capacity than the feature-based and pre-training-based methods. The AUC of Perplexity for the four LLMs is as follows: GPT3.5 (81.64%), GPT-4o (80.88%), Gemini (81.78%), and Llama (79.55%), with a range of 2%. The AUC of pre-training-based methods for each LLM is as follows: GPT3.5 (92.21%), GPT-4o (90.30%), Gemini (75.21%), and Llama (83.26%). The max-min difference is over 17%. These differences can be explained that pre-training-based methods are trained on the code generated by GPT-3.5, while Perplexity is LLM-independent. When it comes to feature-based methods, Perplexity outperforms Idialu and Bukhari. The AUC of Idialu for each LLM is as follows: GPT3.5 (91.43%), GPT-4o (83.16%), Gemini (62.77%), and Llama (71.76%). The max-min difference is 29%. the AUC of Bukhari for each LLM is as follows: GPT3.5 (98.63%), GPT-4o (98.49%), Gemini (92.13%), and Llama (97.15%), with a range of 7%. The max-min range of Bukhari is relatively small. Bukhari only works for C in which HaCode is more different from LLMgCode than other languages, making the detection method easier to detect LLMgCode from HaCode.
$\textbf{RQ}_{\textbf{2}}$: Perplexity has a better generalization capability than the feature-based and pre-training-based methods. The AUC of supervised methods drops significantly when training and testing sets are generated by different LLMs.
V-E $RQ_{\text{3}}$: Inference Efficiency
TABLE IX: The detection speed of Perplexity in different LLMs (seconds per sample)
Model Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
GPT-3.5 Incoder 0.122 0.092 0.111 0.128 0.235 143.5 5.306 5.306
Phi-1 0.148 0.118 0.137 0.151 0.282 170.7 6.307 6.308
StarCoder 0.273 0.231 0.248 0.267 0.508 152.0 11.12 11.12
CodeGen2 0.328 0.281 0.306 0.319 0.606 148.3 13.27 13.28
WizCoder 0.640 0.561 0.577 0.656 1.265 156.1 30.27 30.27
CodeLlama 0.106 0.076 0.096 0.135 0.240 146.7 5.545 5.545
Average 0.270 0.227 0.246 0.276 0.523 152.9 11.97 11.97
GPT-4o Incoder 0.127 0.099 0.118 0.134 0.248 143.1 5.619 5.620
Phi-1 0.157 0.126 0.146 0.163 0.305 143.5 6.738 6.740
StarCoder 0.292 0.249 0.265 0.289 0.553 148.9 11.97 11.97
CodeGen2 0.360 0.303 0.321 0.350 0.662 151.0 14.29 14.29
WizCoder 0.691 0.604 0.623 0.708 1.379 155.2 33.08 33.08
CodeLlama 0.108 0.080 0.100 0.139 0.248 143.9 5.800 5.800
Average 0.289 0.244 0.262 0.297 0.566 147.6 12.92 12.92
Gemini Incoder 0.122 0.094 0.113 0.130 0.238 173.0 5.370 5.370
Phi-1 0.150 0.121 0.140 0.151 0.285 146.1 6.400 6.400
StarCoder 0.278 0.237 0.255 0.267 0.509 169.9 11.14 11.14
CodeGen2 0.336 0.287 0.306 0.328 0.618 146.8 13.37 13.37
WizCoder 0.654 0.575 0.591 0.672 1.288 161.0 30.31 30.31
CodeLlama 0.105 0.078 0.098 0.132 0.237 143.8 5.569 5.570
Average 0.274 0.232 0.251 0.280 0.529 156.8 12.03 12.03
Llama Incoder 0.124 0.096 0.114 0.139 0.248 172.0 5.406 5.407
Phi-1 0.153 0.123 0.141 0.155 0.289 165.7 6.479 6.481
StarCoder 0.283 0.241 0.258 0.273 0.522 162.7 11.38 11.38
CodeGen2 0.342 0.292 0.312 0.328 0.622 150.8 13.58 13.58
WizCoder 0.670 0.586 0.603 0.425 0.801 223.7 16.39 16.39
CodeLlama 0.106 0.079 0.099 0.124 0.217 142.6 4.879 4.879
Average 0.280 0.236 0.255 0.241 0.450 169.6 9.685 9.685
TABLE X: The detection speed of feature-based and pre-training-based methods in different LLMs (seconds per sample)
LLM Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
GPT-3.5 0.038 4.352 0.014
GPT-4o 0.042 4.312 0.010
Gemini 0.040 4.504 0.010
Llama 0.040 4.794 0.010
Table IX and Table X present the inference speed of Perplexity and the other two classes of methods for code generated by four LLMs, respectively. For presentation clarity, we highlight notable results with blue backgrounds (sec/sample:$0-0.01$;$0.01-0.1$;$0.1-1$)
Perplexity is not as fast as the feature-based and pre-training-based methods in all four LLMs. The AUC of Perplexity and the pre-training-based method for each LLM are as follows: GPT3.5 (0.076–0.014), GPT-4o (0.080–0.010), Gemini (0.078–0.010), and Llama (0.079–0.010). Perplexity calculates the perplexity of each token (a code snippet consists of a set of tokens) using another LLM, therefore, it is very time-consuming. When compared with the feature-based method, Perplexity is slower than Idialu: GPT3.5 (0.076–0.038), GPT-4o (0.080–0.042), Gemini (0.078–0.040), and Llama (0.079–0.040). However, Perplexity is faster than Bukhari: GPT3.5 (0.076–4.352), GPT-4o (0.080–4.312), Gemini (0.078–4.504), and Llama (0.079–4.794). Since Bukhari extracts n-gram features from abstract syntactic trees, it has a slow inference speed.
When comparing the inference speed of methods in Perplexity, we observe that methods without perturbation are much faster than methods with perturbation. The inference speed of the worst-performing method without perturbation is less than 2 seconds per sample: GPT3.5 (1.265), GPT-4o (1.379), Gemini (1.288), and Llama (0.801). However, the inference speed of the best-performing method with perturbation is more than 4 seconds per sample: GPT3.5 (5.306), GPT-4o (5.619), Gemini (5.370), and Llama (4.879). Perplexity with perturbations introduces variations into the original sample to generate a list of perturbed samples and calculate the attributes such as log-rank of two types of samples. When comparing Perplexity methods with perturbation, NPR and DetectCodeGPT have a faster inference speed than DetectGPT. The average speed of NPR and DetectCodeGPT is around 10 seconds per sample: GPT3.5 (11.97–11.97), GPT-4o (12.92–12.92), Gemini (12.03–12.03), and Llama (9.685–9.685). NPR and DetectCodeGPT perform very closely because DetectCodeGPT is based on NPR. However, the average inference speed of DetectGPT is more than 100 seconds per sample: GPT3.5 (152.9), GPT-4o (147.6), Gemini (156.8), and Llama (169.6). The perturbation method of DetectGPT is complex, i.e., masking and fulfilling tokens by another LLM. The perturbation method of DetectCodeGPT is simple, i.e., adding spaces and new lines to the original tokens.
$\textbf{RQ}_{\textbf{3}}$: Perplexity has a lower inference speed than the feature-based and pre-training-based methods. The inference speed of Perplexity without perturbation is faster than Perplexity with perturbation.
VI Implications
This section discusses implications based on the experimental results and the detection mechanism of Perplexity.
VI-A For Researchers
Our implications for researchers include both pros and cons of Perplexity.
  • Perplexity is unsuitable for high-level programming languages. Perplexity performs well in low-level languages such as C/C++ (91%) but poorly in high-level languages such as Python (64%) and Ruby (73%), which implies that Perplexity has limited capability for high-level languages. High-level languages encourage a concise and clear code style. High-level languages such as Python provide many built-in operations for basic elements, e.g., list and dictionary; low-level languages have to manually manage array, pointer, etc. The programming patterns of high-level languages can be easier learned by a code generator than that of low-level languages. Moreover, high-level languages have more available third-party libraries than low-level languages. Developers are used to leveraging third-party libraries in their daily development. Since humans and LLMs use third-party libraries in generating code, detecting LLMgCode from HaCode using Perplexity in high-level languages is challenging.
  • Perplexity is unsuitable for short code solutions. Perplexity achieves an AUC of 87.81% for code solutions with more than 50 LOC, but its performance drops to 69.75% for code solutions with less than 20 LOC. This result implies that Perplexity is ineffective for short solutions. Perplexity is based on the probability that an LLM assigns to the next token being the ground-truth token based on previous tokens. Therefore, the perplexity of the first few tokens of code snippets (i.e., a collection of tokens) is unstable due to the lack of contexts. In this case, the perplexity of the first few tokens is relatively high. When the length of code snippets is short, the high perplexity of the first few tokens significantly affects the average perplexity of the entire code snippets. In this case, the perplexities of LLMgCode and HaCode are both high, which makes Perplexity confused.
  • Perplexity is unsuitable for just-in-time detection. NPR and DetectCodeGPT are the best-performing methods in Perplexity. The inference speeds of NPR and DetectCodeGPT are all 10 seconds per sample. On the contrary, the inference speed of the pre-training-based method is less than 0.02 seconds per sample. The difference implies that Perplexity is not suitable for just-in-time detection. This can be explained that Perplexity applies an LLM to calculate the perplexity of each token. The calculation efficiency of LLMs is much lower than small pre-trained models such as CodeBERT. Moreover, NPR and DetectCodeGPT are perturbation-based methods. Such a method calculates the perplexity for one original sample and a collection of perturbed samples. As the number of perturbed samples increases, the perturbation-based methods get time-consuming.
Although our large-scale investigation indicates that Perplexity has several shortcomings, it exhibits advantages in the following aspects.
  • Perplexity exhibits good interpretability. Perplexity tries to simulate code generation mechanisms of LLMs to calculate the perplexity. Therefore, the perplexity value of a code snippet is interpretable. A code snippet with a high perplexity indicates that it is disjointed, which might not be LLMgCode. The threshold of perplexity for detection can be adjusted. Although several supervised methods perform better than Perplexity, their mechanisms are complex and hard to interpret.
  • Perplexity exhibits overall good generalization capability regarding LLMs. Perplexity achieves similar performance in detecting LLMgCode across four LLMs, with AUC variations of around 2%. This implies that Perplexity has a good generalization capability when detecting LLMgCode that is generated by different LLMs. The detection mechanism of Perplexity is similar to LLMs for code generation. Therefore, Perplexity is useful as long as LLMs are based on the auto-regressive generation mechanism (i.e., generating a token based on tokens generated before). The performance of supervised methods drops when the code snippets for training and testing are generated by different LLMs. Considering new powerful LLMs are emerging, Perplexity is the most promising method to address this issue.
  • Perplexity has the potential for fine-grained detection. Perplexity calculates the perplexity of each token in the code snippet. This implies that Perplexity can be applied for fine-grained detection. For instance, if a code snippet includes both LLMgCode and HaCode, Perplexity can calculate the perplexity for each line and detect these lines with low perplexity as LLMgCode. However, feature-based and pre-training-based methods only output a binary result, i.e., the code is either LLMgCode or HaCode.
In short, Perplexity is not effective enough but is the most promising method overall. Unlike those supervised methods that aim to learn characteristics of one specific type of code (e.g., GPT 3.5-generated code), Perplexity seeks to understand intrinsic differences between LLMgCode and HaCode. Therefore, Perplexity has a good generalization capacity. By continuously improving Perplexity on the target dataset, it is expected that the improved performance can be generalized. Moreover, Perplexity can be applied naturally to fine-grained detection (e.g., line level) when LLMgCode and HaCode are mixed, whereas other methods only provide a binary classification result. This enhances the applicability of Perplexity in complex real-world scenarios.
VI-B For Tool Development and Usages
When detecting LLMgCode from HaCode, we should first detect whether it is LLMgCode or whether it is entirely or partially generated by LLMs, then detect which parts (lines or blocks) are generated by LLMs.
The first step can be formulated as a classification problem. However, the experimental results show that none of the three detection methods (classifiers) are effective. Perplexity only performs well for detecting code in C/C++ (from the perspective of language), and large (from the perspective of scale). The feature-based and pre-training-based methods have high effectiveness but poor generalization capability. Overall, they are all acceptable with an AUC of more than 70%. An ensemble learning approach can make full use of each separate method’s advantages meanwhile overcoming disadvantages to some extent. Taking detecting C code as an example, we recommend using DetectCodeGPT (Perplexity), Bukhari (feature-based method), and GPTSniffer (pre-training-based) to develop an ensemble detection tool. By employing an ensemble learning strategy such as MoE (Mixture Of Experts), the detection tool can automatically learn which method(s) to prioritize under different contexts.
The second step is fine-grained detection, i.e., identifying which lines are generated by LLMs. Having understood the code is generated by LLMs, the second step is to first calculate the perplexity of each token in the code and then calculate the average perplexity for each line. These lines with low perplexity are detected as LLMgCode. We recommend highlighting LLMgCode and HaCode in different colors to remind developers.
The detection tool can be applied in but not limited to the following scenarios.
  • Software Supply Chain Management: LLMgCode has become a new third-party asset in software supply chain development [15]. The detection tool helps enterprises manage LLMgCode for tracking accountabilities, checking violations, etc.
  • Code Review: Since LLMgCode is more prone to include defects and vulnerabilities, attention should be placed on it during code review. The detection tool is useful for highlighting LLMgCode to remind reviewers to perform additional checks such as plagiarism detection.
  • Productivity Assessment: Many developers use LLMs to assist programming, making traditional LOC-based productivity assessment less useful. The detection tool helps enterprises assess developers’ productivity more accurately than before.
  • Originality Assessment: LLMgCode is sometimes not considered original. The detection tool is useful for situations where originality is emphasized, including but not limited to software competitions and student assignments.
VII Limitation
This section discusses the possible limitations of this work and our efforts to mitigate their effects.
Test Scale
Due to the budget and time costs, we randomly tested 500 samples for each group in each experiment. For instance, the groups from the perspective of difficulty are easy, middle, and hard. The scale of our testing for each group is comparable to previous related work [14, 16]. Our testing scale is considered acceptable to mitigate biases introduced by random sampling.
Subject Selection
For the experimental subjects, we included several detection methods that were not originally proposed to detect LLMgCode, e.g., DetectGPT and NPR. However, they were proposed to detect LLM-generated text. As source code can be considered as one specific type of text, we believe these detection methods are also suitable for detecting LLMgCode. This setting was also used in previous related work [14, 16]. Moreover, although Perplexity is an unsupervised method, we compared it against supervised methods because Shi et al. [14] reported that Perplexity outperformed supervised methods. We excluded methods whose implementation details are unclear. For example, we excluded AIgCode Detector [26] because it did not specify how perturbations are performed.
Dataset Diversity
For the diversity of HaCode, we collected HaCode from CodeNet only. CodeNet collects HaCode solutions from 14K submits. The HaCode solutions in CodeNet are in 55 languages. 54% of HaCode solutions in CodeNet are accepted, 30% are marked as wrong solutions, and the remaining suffer from one of the possible rejection causes. The richness of CodeNet guarantees the diversity of HaCode in our experiments. For the diversity of LLMgCode, we only evaluated Perplexity in detecting LLMgCode generated by GPT-3.5 in $\text{RQ}_{1}$, which may bias experimental results. However, in $\text{RQ}_{2}$ and $\text{RQ}_{3}$, we found that Perplexity achieved similar results in detecting LLMgCode generated by GPT-4o, Gemini, and Llama. The results show that using code generated by GPT-3.5 is acceptable.
VIII Conclusion
LLMgCode has become increasingly common in software development; however, its security and compliance remain challenging. Moreover, LLMgCode is often mixed with HaCode. Human-LLM code authorship attribution is not only important to improve the transparency of software development, but also to increase the trust of third-party software in the supply chain. In this article, we conduct a large-scale investigation of Perplexity, the state-of-the-art method for detecting LLMgCode from HaCode. Toward realistic evaluations, we use 11,664 HaCode snippets and 13,164 LLMgCode snippets to evaluate Perplexity from three perspectives: detection accuracy, generalization capability, and inference efficiency. The experimental results show that Perplexity is less effective than other methods. Perplexity only performs well in C/C++ from the perspective of language and in large-scale from the perspective of scale. Additionally, Perplexity has a slower inference speed than other methods. However, Perplexity is still the most promising method because it has the best generalization capacity. Furthermore, Perplexity has good interpretability and can be applied for fine-grained detection. In the future, we are going to continually improve Perplexity to detect LLMgCode from HaCode and evaluate it in industry.