### v2

# One Size Does Not Fit All: Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

Abstract.
Large language model-generated code (LLMgCode) has become increasingly common in software development. So far LLMgCode has more quality issues than human-authored code (HaCode). It is common for LLMgCode to mix with HaCode in a code change, while the change is signed by only human developers, without being carefully examined. Many automated methods have been proposed to detect LLMgCode from HaCode, in which the perplexity-based method (Perplexity for short) is the state-of-the-art method. However, the efficacy evaluation of Perplexity has focused on detection accuracy. Yet it is unclear whether Perplexity is good enough in a wider range of realistic evaluation settings. To this end, we carry out a family of experiments to compare Perplexity against feature- and pre-training-based methods from three perspectives: detection accuracy, detection speed, and generalization capability. The experimental results show that Perplexity has the best generalization capability while having limited detection accuracy and detection speed. Based on that, we discuss the strengths and limitations of Perplexity, e.g., Perplexity is unsuitable for high-level programming languages. Finally, we provide recommendations to improve Perplexity and apply it in practice. As the first large-scale investigation on detecting LLMgCode from HaCode, this article provides a wide range of findings for future improvement.
Code authorship attribution, human-authored code, LLM-generated code, large language model, perplexity
journal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Software and its engineering Software verification and validationccs: Software and its engineering Software supply chain security
1. Introduction
The advent of Large Language Models (LLMs) has profoundly advanced automated software engineering (Chen et al., 2021), of which code generation is most popular (Gu et al., 2024). Code generation is to generate source code based on natural descriptions (Li et al., 2023). Various LLMs (e.g., GitHub Copilot (Github, 2025), GPT-4 (OpenAI, 2025)) have been proposed for code generation in recent years (Du et al., 2024). LLMs have the strong capability to follow instructions (Chung et al., 2024), enabling developers to write code by providing natural language descriptions of requirements only (Jiang et al., 2024). Leveraging natural language understanding and generative power, LLMs achieve remarkable success in code generation (Austin et al., 2021). Therefore, LLM-generated code (LLMgCode) has become increasingly common in software development (Sun et al., 2024; Yu et al., 2024a).
Although LLMgCode help developers reduce manual efforts and focus on high-level tasks (Liu et al., 2024a), the quality and security of LLMgCode are challenging (Pearce et al., 2022; Asare et al., 2023; Fu et al., 2023; Liu et al., 2024b). According to Pearce et al. (2022), 40.73% of the 1,689 programs generated by GitHub Copilot are vulnerable, with risks far more than human-authored code (HaCode(Perry et al., 2024; Sandoval et al., 2023). Moreover, LLMgCode has the risks of plagiarism and copyright infringement (Guo et al., 2023), which are less in HaCode. According to Guo et al. (2024), LLMgCode can be a copy of HaCode even if the latter is with copyright protection (Yu et al., 2023). Due to this, GitHub Copilot was accused of compensating \$9 billion for generating restrictive code without attaching the necessary copyright or licenses (Hale, 2022).
In addition to the challenges of LLMgCode itself, the mixture of LLMgCode and HaCode has two challenges. The first is compromising the transparency of software development, which results in inaccurate attribution of code ownership in bug accountability and obscured productivity assessments (Shi et al., 2025). The second is increasing the difficulty of evaluating the reliability of third-party software, which results in distrust for upstream dependencies in the software supply chain (Bukhari et al., 2023). Therefore, It is important to detect LLMgCode from HaCode. To this end, many studies (Shi et al., 2025; Suh et al., 2025) and commercial tools (Copyleaks, 2025) have proposed to detect LLMgCode from HaCode, which supports mitigating various risks by taking measures such as conducting stricter security inspections and performing code plagiarism detection (Xu et al., 2025).
The past few years have seen many studies (Shi et al., 2025; Bukhari et al., 2023; Idialu et al., 2024; Nguyen et al., 2024; Solaiman et al., 2019; Gehrmann et al., 2019; Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023) that proposed methods to detect LLMgCode from HaCode. According to the detection mechanisms these methods use, they can be divided into three categories: (1) feature-based, (2) pre-training-based, and (3) perplexity-based methods (Perplexity for short). Feature-based methods (Bukhari et al., 2023; Idialu et al., 2024) are generally conventional machine learning classifiers that require professionals to devise appropriate features from various aspects such as lexical and syntactic characteristics. Pre-training-based methods (Nguyen et al., 2024) take the code snippets as inputs. They convert code snippets to tokens, embed them as vectors, and use the pre-trained model as a classifier. Perplexity  [^1] For succinctness, Perplexity refers to the family of perplexity-based methods throughout the paper unless it is specifically stated otherwise. (Shi et al., 2025; Solaiman et al., 2019; Gehrmann et al., 2019; Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023) is a class of heuristic methods, which assume that LLMgCode has a lower perplexity than HaCode. Technically, perplexity is a measure of how confident a probability model is in predicting a sample (Xu and Sheng, 2024). When it comes to code generation, perplexity is used to measure the predictability of the next token of code based on prompts and existing tokens. The calculation of the perplexity is based on an LLM.
Figure 1. An example of calculating the perplexity of a set of existing code snippets
Figure 1 presents an example of calculating the perplexity of a probability model (CodeLlama) for code snippets (either LLMgCode or HaCode). The prompt is shown in lines 2-3, and the code snippets to be calculated are in line 6. Based on the prompt, LLM calculates the probability of “  ” (Tab) as the first token is 0.20. Based on the prompt and “  ”, the LLM calculates the probability of “return” as the second token is 0.31. Followed by “ tf” and “.train” are 0.98 and 0.99, respectively. The higher the probability of a token, the more confident the model is about generating this token. Therefore, the perplexity of the model in understanding this token is low. Perplexity is a probability-based measure and has many implementations (e.g., log of probability (Solaiman et al., 2019), log of rank (Ippolito et al., 2019), and log of probability divided by log of rank (Su et al., 2023)) in which the log of probability is common. For example, the perplexity of CodeLlama for “return” is $ln(0.31)=-1.18$. The perplexity of a model for a piece of code is usually the average perplexity of tokens.
When the code is natural and coherent, the model is confident in predicting the next token based on the prompts and the existing tokens, resulting in a lower perplexity. According to the mechanism of LLMs for code generation, LLMs generate code in the unit of a token through understanding previous tokens, calculating the probability of each possible next token, and applying the sampling strategy to select one with high probability as the output. Compared with LLMgCode, HaCode is more casual and personalized, making it more unpredictable. Therefore, it is reasonable that LLMgCode has lower perplexity than HaCode. Due to its explainability and scalability (e.g., log of probability, log of rank), Perplexity has been widely used (Solaiman et al., 2019; Gehrmann et al., 2019; Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023).
Although Perplexity is becoming the most popular method, evaluating Perplexity is limited to detection accuracy. For example, Shi et al. (2025) evaluated the proposed DetectCodeGPT using AUC (Area Under the Curve) on Python only. Yet it is unclear what the detection accuracy of DetectCodeGPT is other than Python. DetectCodeGPT detects only the code generated by LLMs with fewer than 7B parameters; its effectiveness on the commonly used commercial models such as GPT-4o and Gemini-1.0 is unclear. Moreover, DetectCodeGPT uses the same model for code generation and perplexity calculation. As it is hard to know in advance what model is for code generation, the effectiveness of DetectCodeGPT in a wide range of contexts is unclear. The limited evaluation of Perplexity does not reflect its effects well in practice.
Toward realistic evaluations of Perplexity, we conduct a large-scale investigation on Perplexity from three perspectives: detection accuracy, detection speed, and generalization capability. Detection accuracy is the most fundamental goal of detecting LLMgCode, while it might not be good between situations, e.g., an effective method for C++ may not suit Python. In this article, we evaluate the accuracy in various situations from three aspects. Specifically, the detection accuracy (effectiveness) is evaluated from the aspects of the programming language (C, C#, C++, Go, Java, Python, Ruby, and Rust), degree of difficulty (easy, middle, and hard), and scale of solution (small, middle, and large). Efficiency greatly affects the user experience and large-scale applications; thereby, the detection speed (seconds/sample) of Perplexity should be considered. Since new powerful LLMs are emerging, developers utilize various LLMs to assist in programming, the generalization capability of Perplexity is important. The generalization capability is evaluated from the aspect of code generation models (GPT-3.5, GPT-4o, Gemini-1.0, and Llama-3.1). We hope the multi-perspective efficacy evaluations can reflect the efficacy of Perplexity in practice. With experimental evaluation, we first devised a dataset containing 729 programming puzzles, each accompanied by at least two HaCode solutions in each of the eight programming languages available in CodeNet (Puri et al., 2021). To obtain LLMgCode solutions, we used GPT-3.5 to generate two solutions per puzzle in each language, resulting in a total of 11,664 LLMgCode solutions. For each specific evaluation aspect, we randomly sampled 500 LLMgCode and 500 HaCode solutions for each category. For example, when it comes to programming languages, we sampled 500 LLMgCode and 500 HaCode solutions for each language. To evaluate the generalization capability of Perplexity, we generated 500 LLMgCode solutions using each of GPT-4o, Gemini-1.0, and Llama-3.1, ensuring a consistent sample size with other evaluation aspects. Specifically, we randomly selected 32 puzzles from the dataset and used each LLM (GPT-4o, Gemini-1.0, and Llama-3.1) to generate two solutions per puzzle in each language. Each LLM thus produced 512 LLMgCode solutions, from which we randomly sampled 500 to form the final set of its solutions. Then, we compared eight perplexity-based methods against two feature-based methods and one pre-training-based method that have been used to detect LLMgCode from HaCode.
The experimental results show that (1) for detection accuracy, the best-performing Perplexity performs well in detecting LLMgCode in C/C++ (AUC=91%) and large-scale code (AUC=88%) but still does not perform as well as the other two methods regarding programming language, degree of difficulty, and scale of solution; (2) for detection speed, Perplexity without perturbation is as efficient as others, while Perplexity with perturbation have a relatively slow efficiency, with detection speed ranging from 10 to 150 seconds per sample; (3) for generalization capability, Perplexity perform better than others across different LLMs, with AUC ranging from 80% to 82%. Based on experimental results, we provide implications of Perplexity from both positive and negative perspectives. On the negative side, Perplexity is unsuitable for high-level programming languages, short code snippets, and just-in-time detection. On the positive side, Perplexity exhibits good generalization capability, which performs well for C/C++, and has good interpretability. Finally, we provide three directions for improving Perplexity in the future and recommendations for the application of Perplexity.
The main contributions of this article can be summarized as follows.
  • Experimental design: An experimental design to investigate the detection accuracy, detection speed, and generation capability of Perplexity in detecting LLMgCode.
  • Findings: A wide range of findings that illustrate the advantages and limitations of Perplexity in detecting LLMgCode.
  • Dataset: A large-scale dataset for continuous research on detecting LLMgCode.
The remainder of this article is organized as follows. Section 2 presents background and related work. Section 3 illustrates motivating examples. Section 4 and Section 5 elaborate on the research methodology and results, respectively. Section 6 discusses implications. Finally, we present threats to validity in Section 7 and conclude this article in Section 8.
2. Background and Related Work
This section presents an overview of LLMs for code generation and reviews studies that apply Perplexity for code authorship attribution.
2.1. Background: Large Language Models for Code Generation
Since LLMs were proposed, they have been widely applied in many tasks across a variety of domains (Bommasani et al., 2021; Wei et al., 2022; Zhao et al., 2023; Wu et al., 2023). One of the most exciting applications of LLMs is code generation (Brown et al., 2020; Dong et al., 2024; Hong et al., 2024; Zhang et al., 2023; Zhou et al., 2023). LLMs are capable of generating source code based on prompts. The prompts consist of natural and programmatic descriptions of code functions, retractions, input and output formats, or any other elements. Following the prompts, an LLM generates tokens sequentially until it arrives at a stop indicator or the token limit.
Many studies show that LLMgCode helps developers save time (Coignion et al., 2024), reduce effort (Ziegler et al., 2022), and improve productivity (Siddiq et al., 2024) in software development. Moreover, LLMgCode help ease developers as they sometimes suffer from heavy workload (Dunne et al., 2024). With LLMgCode, developers can pay more attention to important and urgent tasks. Because of these benefits, LLMgCode has been widely used and has even become a must-have in modern software development (Spiess et al., 2025; North et al., 2024). A survey made with 500 USA developers showed that 92% of respondents use LLMgCode for official assignments and personal use (Shani and Staff, 2023).
Although LLMs bring many benefits, the quality and security of LLMgCode remain challenging (Pearce et al., 2022; Yu et al., 2024b; Dakhel et al., 2023). One of the causes is that the corpus used for training LLMs is not qualified and secure enough (Asare et al., 2023). Therefore, LLMgCode might be buggy and vulnerable (Schuster et al., 2021; Jesse et al., 2023). For example, Fu et al. (2022) assessed the security of LLMgCode in six common programming languages and found that 156 out of 435 code segments contained security vulnerabilities. The proportion of vulnerabilities in C++ and Go code accounts for 46.1% and 45.0%, respectively. Mastropaolo et al. (2023) assessed the robustness of LLMgCode and found that slight modifications to the natural description while still maintaining semantic equivalence would result in LLMs generating different results, occurring in 46% (408/892) of the total samples. Jimenez et al. (2023) assessed the capability of large language models to solve 2294 real-world problems (Issues) on GitHub and found that the pass rate of code generated using GPT-4 has the lowest pass rate–0%, using Claude 2 has the highest pass rate–1.96%. In addition, LLMgCode might be similar or even unchanged from existing code snippets in the training corpus (Guo et al., 2024). As the training corpus might include a significant amount of restrictive code (Yu et al., 2023), there are copyright infringement risks for LLMgCode.
2.2. Related Work: Perplexity for Code Authorship Attribution
Detecting LLMgCode from HaCode can be regarded as a code authorship attribution problem (Abuhamad et al., 2018; Dauber et al., 2018; Bogomolov et al., 2021) in which the authors are either LLMs or humans. Code authorship attribution fills in the missing or inaccurate author information of source code fragments, which is beneficial for software maintenance (Girba et al., 2005; Anvik et al., 2006; Fritz et al., 2010) and software quality analysis (Rahman and Devanbu, 2011; Bird et al., 2011; Yin et al., 2011; Thongtanunam et al., 2016). For example, code authorship attribution helps detect code plagiarism (Burrows and Tahaghoghi, 2007; Kothari et al., 2007) and trace malware provenance (Caliskan-Islam et al., 2015). Moreover, detecting LLMgCode from HaCode increases the authenticity and integrity of software (Shi et al., 2025), reducing a variety of risks such as security vulnerabilities and copyright infringement in the software supply chain (Bukhari et al., 2023). In addition, LLMgCode is typically not considered original (Nguyen et al., 2024), thus detecting LLMgCode is helpful in scenarios that require individual work, such as programming contests (Idialu et al., 2024; Nguyen et al., 2024). This scenario also asks for a high detection efficiency, as examination systems need to provide quick feedback after submission.
Perplexity is originally one of the major methods used for detecting AI-generated content (including but not limited to LLMgCode) (Ippolito et al., 2019; Mitchell et al., 2023; Su et al., 2023). For example, Mitchell et al. (2023) proposed DetectGPT to detect LLM-generated text. Technically, DetectGPT utilizes the log probabilities to calculate the perplexity. The authors evaluated DetectGPT using the AUC metric on 500 LLM-generated samples and 500 human-written samples. In addition, they evaluated the generalization capability of DetectGPT by alternately using GPT-J-6B, GPT-Neo-2.7B, and GPT-2-1.5B as the code generation model and the perplexity calculation model. Su et al. (2023) proposed two methods to calculate the perplexity. The first is Log-Likelihood Log-Rank Ratio (LRR) and the second is Normalized Perturbed Log-Rank (NPR). The authors evaluated LRR and NPR using the AUC metric on 300 LLM-generated and 300 human-written samples. They also compared LRR and NPR against baselines in terms of detection speed.
With the increasing popularity of LLMgCode, Perplexity has been commonly used in detecting LLMgCode from HaCode (Shi et al., 2025; Xu and Sheng, 2024). Xu and Sheng (2024) proposed a method using a variant of probability to calculate the perplexity. They evaluated their methods using the AUC, FPR, and TPR metrics on 5,214 LLM-generated and 5,214 human-authored code snippets. They also evaluated the influence of models used to calculate perplexity. Shi et al. (2025) proposed a method namely DetectCodeGPT to detect LLMgCode from HaCode. DetectCodeGPT calculates the perplexity based on NPR. They evaluated DetectCodeGPT using the AUC metric on 500 LLM-generated and 500 human-authored code snippets. They also evaluated the generalization capability of DetectCodeGPT by alternately using Incoder-1B, Phi-1B, StarCoder-3B, WiardCoder-3B, CodeGen2-3.7B, and GodeLlama-7B as the code generation model and the perplexity calculation model. These studies focus on proposing new perplexity-based detection methods and evaluating overall detection accuracy. As a result, they only partially consider other detection methods and pay limited attention to broader evaluation aspects, leading to limitations in both evaluation objectives and dimensions. To address this gap, this study conducts a systematic evaluation of feature-based, pre-training-based, and perplexity-based methods from multiple perspectives, including detection accuracy (language, difficulty, scale), detection speed, and generalization capability.
Overall, Perplexity is a popular method for detecting LLMgCode from HaCode but its evaluation is limited. In this article, we conduct a large-scale investigation on the efficacy of Perplexity from multiple aspects.
3. Motivating Examples
This section describes three examples to motivate this investigation.
3.1. Code Authenticity
Although Perplexity has been demonstrated useful in detecting LLMgCode, the LLMgCode used in previous work is generated by “tiny” LLMs, such as CodeGen2-3.7B (Shi et al., 2025). Their effectiveness for larger code generators is unknown.
Figure 2. An example of the LLMgCode (CodeGen2-3.7B) and HaCode
Figure 2 presents examples of LLMgCode and HaCode. The prompt indicates that the function to be generated is to synchronize the file state across editor stacks to prevent unnecessary reloads after a save. The prompt is the input of CodeGen2-3.7B (shown in the first block). The LLMgCode is the solution generated by CodeGen2-3.7B (shown in the second block), and the HaCode is the solution provided by a human developer (shown in the third block). The LLMgCode is an incorrect implementation of this functionality and differs significantly from HaCode. In fact, the LLMgCode only returns when the file was not saved under a new name, which is inconsistent with the intended functionality. While HaCode avoids unnecessary reloads by updating the file state: it locates the original filename and synchronizes its metadata, including the creation status, filename, and newest modified time. Although it is easy to detect such LLMgCode from HaCode, few developers use such LLMgCode for their code solutions, as it is far from addressing real-world problems.
When Perplexity uses the same model for code generation and detection, it achieves low perplexities of LLMgCode. This is because the model would assign a high prediction probability to each token when it perceives the code generated by itself. However, in real-world scenarios, the detection model does not have clues to know which model is used to generate code early. Moreover, these LLMs used for generating code have hundreds of billions of parameters, making them challenging to deploy locally to calculate perplexities.
3.2. Language Diversity
Perplexity has been evaluated in limited programming languages so far. However, the code stylometry of different programming languages varies greatly. Figure 3 presents two pieces of code generated by GPT-3.5, aiming to calculate the median of all subarrays of a given array and then calculate the median of these medians.
Figure 3. An example of code achieving the same task in Python and C
As illustrated in Figure 3, the code stylometries differ in Python and C, particularly in their logic and scale. Python has rich built-in libraries. For example, “statistics.comedian” is a tool library in Python. It can directly calculate the median without the need for manual sorting and selection. While C does not have a direct median calculation function, therefore, we use “qsort” to sort and manually calculate the median. In fact, many high-level functions must be manually implemented in C, such as inserting or deleting elements in arrays. However, such functions can be automatically handled by the built-in methods of lists in Python. In addition, Python is a dynamically typed programming language that does not require declaring variable types, making the code more concise. While C is a statically typed programming language where each variable type needs to be declared. Moreover, manipulating elements such as list slicing is easy in Python. For example, “a[i: j+1]” returns subarrays without manual calculation and traversal. While C requires manually managing memory and arrays, which means that developers need to handle each element carefully.
Due to differences such as the built-in library support and design principles, code written in different programming languages is distinguishable. In addition to programming languages, the code used to solve puzzles of different difficulty and the code of different scales are distinguishable in terms of logic and complexity.
3.3. Generalization Capability
There are many LLMs available for generating code, such as GPT, Copilot, and Gemini. Due to differences in training data, model architecture, optimization objectives, generation strategies, etc., the code generated by these LLMs is different. Figure 4 presents two Java code snippets generated by GPT-4o and Gemini-1.0 solving the same programming puzzle. The puzzle is a dynamic programming problem that is similar to the unbounded knapsack problem. This problem can be described as follows. Given the monster’s initial health $H$ and $N$ different spells, each spell has a fixed damage $A_{i}$ and a magic point cost $B_{i}$. The objective is to reduce the monster’s health to 0 or below by casting spells while minimizing the total magic points consumed.
Figure 4. An example of code achieving the same task generated by GPT-4o and Gemini-1.0
As shown in Figure 4, although both pieces of code aim to reduce the monster’s health to 0 with minimal loss, the two LLMs have different implementations. GPT-4o starts to generate code from the point of view of damage, with the aim of achieving total damage $H$ at the minimum cost. Gemini-1.0 starts to generate code from the perspective of the monster’s health, aiming to reduce the monster’s health to 0 with the minimum cost. The former is to intuitively accumulate damage in the forward direction, and the latter is to reduce the monster’s health in the reverse direction.
In addition to the differences mentioned above, the abstraction levels are different. The solution generated by Gemini-1.0 is detailed and complex. Gemini-1.0 uses a two-dimensional array ‘dp’ to simultaneously record health values and spell combinations. GPT-4o uses only a one-bit array ‘dp’ to record the minimum cost of achieving each damage value. When updating the state, Gemini-1.0 considers all spell combinations in each update, making it more complex, while GPT-4o directly updates the state by traversing damage values and spell methods. In addition, Gemini-1.0 considers boundary conditions and special cases, such as returning an outlier of “-1” when the target cannot be reached. Unfortunately, GPT-4o fails to handle these boundary conditions, reducing its robustness.
In short, GPT-4o focuses on core logic, while Gemini-1.0 focuses on practical usages, boundary conditions, and robustness regarding this example. This example motivates us to investigate the generalization capability of Perplexity when the code is generated by different LLMs. In addition to generalization, the detection speed of Perplexity is important, as it influences user experience.
4. Research Methodology
This section elaborates on the experimental designs to investigate the efficacy of Perplexity in detecting LLMgCode.
4.1. Research Questions
Existing work focuses on evaluating the detection accuracy of Perplexity, however, there are many aspects to consider when using it in practice. In this article, we investigate the efficacy of Perplexity in detecting LLMgCode in a wider range of realistic evaluation settings.
  • $\text{{RQ}}_{\textbf{1}}$:
    Is Perplexity an effective method in detecting LLMgCode?: Language ($\text{{RQ}}_{\textbf{1.1}}$), Difficulty ($\text{{RQ}}_{\textbf{1.2}}$), and Scale ($\text{{RQ}}_{\textbf{1.3}}$)
    Motivation:
    Detection accuracy is crucial in detecting LLMgCode, while it might not be consistently good between situations, e.g., an effective method for C++ might not be suitable for Python. We propose $\text{RQ}_{1}$ to compare Perplexity against state-of-the-art methods regarding detection accuracy from three sub-aspects: (1) programming languageof the implementation: C, C#, C++, Go, Java, Python, Ruby, and Rust; (2) degree of difficultyof the puzzle: easy, medium, and hard; and (3) scale of solutionto the puzzle: small, middle, and large.
  • $\text{{RQ}}_{\textbf{2}}$:
    Is Perplexity an efficient method in detecting LLMgCode?
    Motivation:
    Detection speed affects user experience in just-in-time scenarios such as time-limited programming contests, and is important for large-scale applications such as batch scanning of software. According to the detection mechanism, Perplexity applies perturbation strategies to generate dozens of perturbation samples and uses an LLM to calculate the probability or rank of each token for both the original sample and perturbation samples (Shi et al., 2025; Su et al., 2023). Such a two-phase detection process might be time-consuming. In $\text{RQ}_{2}$, we compare Perplexity with feature- and pre-training-based methods regarding detection speed.
  • $\text{{RQ}}_{\textbf{3}}$:
    Does Perplexity have the generalization capability?
    Motivation:
    Developers utilize various LLMs to assist in programming. Due to differences in training data, model architecture, and generation strategies, the code generated by different LLMs varies significantly. An optimal detection method should be capable of detecting the code generated by different LLMs. In $\text{RQ}_{3}$, we investigate the generalization capability of Perplexity in detecting LLMgCode.
4.2. Data Preparation
Although there are datasets (Shi et al., 2025; Bukhari et al., 2023; Idialu et al., 2024) that work to detect LLMgCode from HaCode, none of them is suitable to address all research questions. These datasets have the following limitations: (1) they support only Python and C (limited diversity in languages); (2) their LLMgCode code is generated from a single LLM or a single family of LLMs (limited capability in generalization); and (3) only one dataset (Idialu et al., 2024) provides difficulty scores of puzzles, (limited degrees in difficulty). Therefore, we developed a new dataset based on CodeNet (Puri et al., 2021). CodeNet was developed by IBM to promote AI technologies in the code domain and has been widely used for many studies (Xu and Sheng, 2024; Agarwal et al., 2021; Goel et al., 2022; Pan et al., 2024; Naik et al., 2022, 2023). CodeNet comprises a large number of submissions from real-world developers. More than half of these submissions are functionally correct (accepted), and more than 90% are successfully compiled and executed (executable). The high quality of CodeNet ensures our HaCode serves as a reliable baseline.
Figure 5. An example of a programming puzzle in CodeNet
As shown in Figure 5, a programming puzzle provided by CodeNet has seven basic elements: (1) difficulty score, (2) statement, (3) constraint, (4) input, (5) output, (6) sample input, and (7) sample output. The difficulty score measures the difficulty level of the puzzle. The statement describes what the puzzle is. The constraint describes the specific constraints on the variables. The input and output describe the requirements for input and output, respectively. Finally, the sample input and sample output provide the input and output of a specific example, respectively. Based on CodeNet, we developed our dataset through the following five steps.
Figure 6. The prompts used for generating LLMgCode solutions for programming puzzles
  • Step-1: Problem selection. To build an appropriate dataset for evaluation, we excluded the programming problem whose (1) statement is not in English as it is the most popular language; (2) statement contains images or tables as they can hardly be extracted and input into LLMs; (3) solutions are all less than 30 LOC (Lines Of Code) as they are mainly for trivial puzzles; (4) solutions are incomplete regarding all the eight programming languages as they constitute the HaCode to be evaluated in $\text{RQ}_{1.1}$; and (5) structure is not complete, i.e., statement, input, output, constraints, difficulty, and samples ($\geq 2$). After employing these criteria, our dataset includes 729 programming puzzles.
  • Step-2: HaCode collection. For each puzzle selected in Step-1, we randomly took two code solutions for each of the eight programming languages from CodeNet. The code solutions were submitted by human developers, including correct, incorrect, and rejected solutions. Finally, we had 16 HaCode solutions for each puzzle.
  • Step-3: LLMgCode generation. We generated LLMgCode for programming puzzles using OpenAI GPT-3.5 (gpt-3.5-turbo-0125). The prompts for code generation are shown in Figure 6. There are two solutions for each puzzle regarding each of the eight programming languages. Moreover, to address $\text{RQ}_{3}$, we used the prompts from Figure 6 and input them into GPT-4o, Gemini-1.0, and Llama-3.1, with each model generating 500 solutions. Note that the following two steps are performed on the code solutions that have been generated by GPT-3.5.
  • Step-4: Classification according to the difficulty of puzzles. We categorized the puzzles based on their difficulty degrees and set two thresholds for the difficulty score to ensure that each group has the same number of puzzles. The detailed settings are: easy for $\leq 200$, medium for $>200$ and $\leq 500$, and hard for $>500$.
  • Step-5: Classification according to the scale of solutions. We categorized the code solutions based on their scale (LOC) and set two thresholds to ensure that each group has the same number of solutions. The detailed settings are: small for $\leq 20$, medium for $>20$ and $\leq 50$, and large for $>50$.
Finally, we obtained a dataset containing function-level LLMgCode and HaCode that solve the same programming puzzles, which is used for the experimental evaluations. We perform the evaluation on function-level code because (1) due to the current limitations of LLMgCode in handling complex programming tasks, developers usually use it to assist in implementing functions rather than entire software repositories. As a result, repo-level detection also needs to be performed at the function level, and (2) to the best of our knowledge, existing LLMgCode detection methods are all designed for function-level detection. Therefore, we evaluate them using the function-level code to ensure a fair evaluation.
4.3. Studied Methods
We compared Perplexity against feature- and pre-training-based methods.
4.3.1. Feature-based Methods
Feature-based methods require professionals to devise appropriate features of code, e.g., lexical and syntactic features, and input these features into machine learning classifiers for detection. We investigate two feature-based methods as follows.
  • Idialu et al. (Idialu et al., 2024) takes a combination of layout, syntactic, and lexical features of Python programs as the input, where layout features refer to code formatting such as indentation and spacing, lexical features are derived from code tokens such as keywords and literals, and syntactic features capture the structural patterns and relationships between code elements. Idialu uses XGBoost to construct the detection method.
  • Bukhari et al. (Bukhari et al., 2023) takes a combination of lexical and syntactic features of C programs as the input. The lexical features are derived from the feature set introduced by Caliskan-Islam (Caliskan-Islam et al., 2015) for code stylometry. The syntactic features are extracted from abstract syntax trees. Bukhari uses Random Forest, SVM, KNN, and XGBoost to build the detection methods. XGBoost outperforms the other classifiers in terms of detection performance.
4.3.2. Pre-Training-based Method
Pre-training-based methods take code snippets as input, convert them into tokens, embed them as vectors, and leverage a pre-trained model for classification. We investigate one pre-training-based method as follows.
  • GPTSniffer (Nguyen et al., 2024) designs rewriting rules (e.g., rewriting class names) to create diverse derivations of the original code samples. These derivations are then used to fine-tune a pre-trained model, i.e., CodeBERT, as the detector. To perform detection, GPTSniffer transforms the input code into a sequence of independent tokens, which are then fed into the fine-tuned CodeBERT to predict whether the code is LLMgCode.
4.3.3. Perplexity
Eight popular perplexity-based methods are as follows. The first five calculate the perplexity based on the original sample without perturbations. The last three require applying other LLMs to perturb the original sample and calculate the perplexity based on the discrepancies before and after the perturbation.
  • Log-p(x) (Solaiman et al., 2019) utilizes the average token-wise log probability to calculate the perplexity. The log probability is the logarithm of the probability of the next token (ground truth) based on previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a larger log-p(x) than HaCode. The calculation of log-p(x) is shown in Equation 1, where $p_{\theta}(x_{i}\mid x_{<i})$ is the probability of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
[…]
  • Entropy (Gehrmann et al., 2019) uses the average entropy to calculate the perplexity. Entropy is the sum of the probability multiplied by the negative logarithm of the probability of each token. This method assumes that LLMgCode has a larger entropy than HaCode. The calculation of entropy is shown in Equation 2, where $p_{\theta}(x_{i}\mid x_{<i})$ is the probability of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
    (2) $\textrm{Entropy}=-\frac{1}{t}\sum_{i=1}^{t}p_{\theta}(x_{i}\mid x_{<i})\log p_{\theta}(x_{i}\mid x_{<i})$
  • Log-Rank/Rank (Gehrmann et al., 2019; Ippolito et al., 2019) utilizes the average log-rank/rank of each token to calculate the perplexity. Log-Rank/Rank is the logarithm of rank/rank of the next token in the vocabulary list. The vocabulary list is sorted according to the probability of each token given the previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a lower log-rank/rank than HaCode. The calculations of Log-Rank and Rank are shown in Equation 3 and Equation 4 respectively, where $r_{\theta}(x_{i}\mid x_{<i})$ is the rank of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
[…]
    (4) $\textrm{Rank}=\frac{1}{t}\sum_{i=1}^{t}r_{\theta}(x_{i}\mid x_{<i})$
  • LRR (Su et al., 2023) divides log-likelihood by log-rank to calculate the perplexity. Log-likelihood is the average token-wise log probability, representing the absolute confidence for the true token. Log-rank is the average token-wise log rank, representing the relative confidence for the true token. For LLMgCode, log-rank is more discernible than the log-likelihood. This method assumes that LLMgCode has a larger LRR than HaCode. The calculation of LRR is shown in Equation 5, where $p_{\theta}(x_{i}\mid x_{<i})$ is the probability of token $x_{i}$ conditioned on the previous tokens, $r_{\theta}(x_{i}\mid x_{<i})$ is the rank of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
    (5) $\textrm{LRR}=-\frac{\sum_{i=1}^{t}\log p_{\theta}(x_{i}\mid x_{<i})}{\sum_{i=1}^{t}\log r_{\theta}(x_{i}\mid x_{<i})}$
  • DetectGPT (Mitchell et al., 2023) leverages the discrepancy between the log probability of the original sample (a collection of tokens) and the average log probability of the perturbed samples to calculate the perplexity, normalized by the standard deviation of the log probability of the perturbed samples. The perturbed samples are obtained by masking and completing some tokens. This method assumes that LLMgCode has a greater perturbation discrepancy as it is more sensitive to perturbations than HaCode. The calculation of DetectGPT is shown in Equation 6, where $x$ is the original sample, $\tilde{x}$ is the perturbed sample, $\log p_{\theta}(x)$ is the log probability of the original sample $x$, $\log p_{\theta}(\tilde{x}_{i})$ is the log probability of the perturbed sample $\tilde{x}_{i}$, $k$ is the number of perturbed samples.
    (6) $\textrm{DetectGPT}=\frac{\log p_{\theta}(x)-\frac{1}{k}\sum_{i=1}^{k}\log p_{\theta}(\tilde{x_{i}})}{\sqrt{\frac{1}{k-1}\sum_{p=1}^{k}{(\log p_{\theta}(\tilde{x}_{p})-\frac{1}{k}\sum_{i=1}^{k}\log p_{\theta}(\tilde{x}_{i}))}^{2}}}$
  • NPR (Su et al., 2023) divides the average log-rank of the perturbed samples by the log-rank of the original sample to calculate the perplexity. The perturbed samples are obtained by replacing some tokens. This method assumes that LLMgCode has a larger NPR as it is more susceptible to perturbations than HaCode. The calculation of NPR is shown in Equation 7, where $x$ is the original sample, $\tilde{x}$ is the perturbed sample, $\log r_{\theta}(x)$ is the log rank of the original sample $x$, $\log r_{\theta}(\tilde{x}_{p})$ is the log rank of the perturbed sample $\tilde{x}_{p}$, $k$ is the number of perturbed samples.
    (7) $\textrm{NPR}=\frac{\frac{1}{k}\sum_{p=1}^{k}\log r_{\theta}(\tilde{x}_{p})}{\log r_{\theta}(x)}$
  • DetectCodeGPT (Shi et al., 2025) utilizes the discrepancy between the NPR of the original sample and the average NPR of the perturbed samples to calculate the perplexity. The perturbed samples are obtained by adding spaces and new lines to the original sample. This method assumes that the average NPR of perturbed LLMgCode drops more than HaCode. The calculation of NPR is shown in Equation 8, where $x$ is the original sample, $\tilde{x}$ is the perturbed sample, $\log r_{\theta}(x)$ is the log rank of the original sample $x$, $\log r_{\theta}(\tilde{x}_{i})$ is the log rank of the perturbed sample $\tilde{x}_{i}$, $k$ is the number of perturbed samples.
    (8) $\textrm{DetectCodeGPT}=\frac{\frac{1}{k}\sum_{p=1}^{k}\log r_{\theta}(\tilde{x}_{p})}{\log r_{\theta}(x)}-\frac{1}{k}\sum_{i=1}^{k}\frac{\frac{1}{k}\sum_{p=1}^{k}\log r_{\theta}(\tilde{x}_{p})}{\log r_{\theta}(\tilde{x}_{i})}$
When selecting experimental subjects, we included several detection methods that were not originally proposed to detect LLMgCode, e.g., DetectGPT and NPR. They were proposed to detect LLM-generated text. Since source code is one type of text, these detection methods are also suitable for detecting LLMgCode. This setting was also used in previous related work (Shi et al., 2025; Suh et al., 2025). Although Perplexity is an unsupervised method, we compared it against supervised methods because Shi et al. (2025) reported that Perplexity outperformed supervised methods. We excluded methods whose implementation details are unclear. For example, we excluded AIgCode Detector (Xu and Sheng, 2024) because it did not specify how perturbations are performed.
4.4. Experimental Settings
Following related work (Shi et al., 2025; Idialu et al., 2024; Mitchell et al., 2023; Su et al., 2023), we use the AUC (Area Under the Curve) to evaluate each detection method. AUC stands for the area under ROC (Receiver Operating Characteristic curve). Given a set of TPR (True Positive Rates) and FPR (False Positive Rates) across different thresholds, the calculation of AUC is as follows.
[…]
where $t$ denotes the varying threshold values.
For $\text{RQ}_{1.1}$, we evaluated Perplexity and pre-training-based methods in all eight programming languages. We evaluated feature-based methods on specific languages because the features are designed for those languages only. Due to the time and cost limit, we tested 500 pairs of human-LLM code snippets for each aspect regarding each research question. For $\text{RQ}_{1}$, we tested 500 pairs of human-LLM code snippets for each programming language, difficulty level, and scale. For $\text{RQ}_{2}$ and $\text{RQ}_{3}$, we tested 500 pairs of human-LLM code snippets for each LLM. We randomly selected 500 pairs of human-LLM code snippets from the dataset each time. We selected 500 pairs for the reason of budget and time costs, and such a scale is comparable to the prior related work Shi et al. (2025) (500) and Suh et al. (2025) (400). Following Shi et al. (2025), we utilized CodeT5-770M to generate perturbed code snippets for DetectGPT. We added spaces and empty lines to generate perturbed code snippets for NPR and DetectCodeGPT. We generated 50 perturbed samples for each code sample. Code-specific LLMs are pre-trained using code corpora, which equips them with a deeper understanding of programming languages. We use code-specific LLMs to compute the perplexity of code snippets, as they are more capable of understanding code than general LLMs. Following Shi et al. (2025), we utilized Incoder (1.3B), Phi-1 (1.3B), StarCoder (3B), WizardCoder (3B), CodeGen2 (3.7B), and CodeLlama (7B) to calculate the perplexity for each Perplexity. Table 1 summarizes the specific usage of all LLMs involved in our study. We carried out the experiments using 1 NVIDIA A100-PCIE-40GB.
Table 1. Summary of LLM usage
LLM Usage
GPT3.5 To generate 11,664 LLMgCode solutions for all the evaluation aspects.
GPT-4o To generate 500 LLMgCode solutions for the generalization capability.
Gemini-1.0
Llama-3.1
Incoder (1.3B) To calculate perplexity values for all methods in Perplexity.
Phi-1 (1.3B)
StarCoder (3B)
CodeGen2 (3B)
WizCoder (3.7B)
CodeLlama (7B)
CodeT5 (770M) To generate perturbed code snippets for DetectGPT.
5. Results and Analysis
This section reports on results and analysis to answer research questions.
5.1. $\text{RQ}_{\text{1.1}}$: Programming Language
LABEL:tab:performance-perplexity and Table 3 show the AUC of the three groups of detection methods from the perspective of programming languages. The code LLMs listed in LABEL:tab:performance-perplexity are used to calculate perplexity values. Since feature-based detection methods are designed for specific languages, we provide results only in Python (Idialu) and C (Bukhari) in Table 3. The notable results are highlighted in blue (AUC:$90\%-100\%$;$80\%-90\%$;$70\%-80\%$).
Table 2. The AUC of Perplexity in terms of programming languages
Language Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
C Incoder 56.43 38.55 19.07 59.98 64.80 57.31 84.98 84.80
Phi-1 83.66 36.19 59.41 77.13 25.84 61.66 90.17 90.40
StarCoder 65.43 48.50 49.85 66.72 58.72 56.32 86.47 85.70
CodeGen2 74.30 39.72 52.23 72.12 53.33 56.40 88.51 87.08
WizCoder 80.87 30.02 58.61 77.82 53.16 58.65 88.73 89.09
CodeLlama 74.75 41.99 57.49 70.25 58.83 57.27 85.05 84.95
Average 72.57 39.16 49.44 70.67 52.45 57.94 87.32 87.00
C# Incoder 48.75 48.97 53.59 56.97 82.05 33.17 64.66 65.12
Phi-1 90.12 41.50 84.61 86.90 36.59 39.28 84.54 86.11
StarCoder 62.09 55.50 72.55 67.69 71.59 27.34 73.39 73.62
CodeGen2 75.45 45.08 77.50 77.80 68.86 35.67 78.85 78.55
WizCoder 76.73 36.49 75.41 76.48 54.82 33.22 76.45 79.04
CodeLlama 62.96 59.45 72.53 67.66 71.32 35.06 66.94 68.34
Average 69.35 47.83 72.70 72.25 64.21 33.96 74.14 75.13
C++ Incoder 63.99 34.96 19.79 67.89 69.92 56.17 81.90 82.38
Phi-1 88.38 35.20 63.85 82.41 34.02 59.22 91.14 92.10
StarCoder 75.17 44.86 56.75 74.79 70.28 54.83 89.56 89.35
CodeGen2 79.91 39.33 59.04 74.43 57.40 56.00 86.33 86.04
WizCoder 83.06 31.19 61.39 83.11 56.86 57.46 90.52 91.37
CodeLlama 76.00 48.37 59.64 77.39 67.68 56.51 86.08 86.65
Average 77.75 38.99 53.41 76.67 59.36 56.70 87.59 87.98
Go Incoder 41.91 53.58 22.65 43.97 71.24 29.42 54.83 54.47
Phi-1 72.68 45.99 61.24 68.27 40.24 37.04 68.74 70.10
StarCoder 50.95 60.61 52.13 49.62 47.88 32.19 53.33 53.59
CodeGen2 53.17 62.07 54.42 53.88 45.76 35.64 55.68 55.54
WizCoder 64.71 60.73 48.72 57.91 45.15 32.09 62.42 64.21
CodeLlama 47.84 65.36 53.63 49.98 49.25 32.54 53.00 54.40
Average 55.21 58.06 48.80 53.94 49.92 33.15 58.00 58.72
Java Incoder 54.61 44.73 21.18 61.72 67.74 42.03 75.11 74.65
Phi-1 85.44 40.40 62.55 79.07 28.80 49.80 84.69 85.54
StarCoder 63.01 56.90 55.61 61.27 53.33 42.25 74.87 73.67
CodeGen2 71.96 50.69 51.89 66.52 49.20 44.81 72.75 71.94
WizCoder 69.97 45.14 59.85 65.34 47.12 43.20 69.29 70.36
CodeLlama 64.27 57.67 54.88 61.21 56.86 41.67 65.65 65.95
Average 68.21 49.26 50.99 65.85 50.51 43.96 73.73 75.23
Python Incoder 47.33 51.23 44.73 50.34 57.62 41.02 60.91 58.52
Phi-1 60.03 53.48 48.46 56.93 34.82 45.08 63.89 62.64
StarCoder 55.72 53.83 40.02 52.48 37.86 40.96 58.03 55.70
CodeGen2 56.54 53.36 42.20 50.80 34.18 38.70 59.02 56.48
WizCoder 62.79 44.41 47.66 56.25 23.82 42.54 63.61 62.52
CodeLlama 57.11 48.76 43.67 48.84 32.16 40.25 55.94 54.60
Average 56.59 50.84 44.46 52.61 36.74 41.43 60.23 58.41
Ruby Incoder 52.41 44.31 46.83 51.24 61.11 39.17 55.49 54.16
Phi-1 67.47 39.48 63.86 65.58 54.34 44.45 66.32 65.64
StarCoder 61.61 44.25 47.51 59.77 42.90 38.70 60.04 57.52
CodeGen2 63.78 40.34 53.65 59.29 45.46 41.44 63.95 61.61
WizCoder 70.22 33.63 50.58 66.85 41.83 44.25 73.44 71.65
CodeLlama 61.63 42.48 57.81 65.62 58.55 39.15 73.17 72.05
Average 62.85 40.75 53.37 61.39 50.70 41.19 65.40 63.77
Rust Incoder 68.41 38.70 69.87 71.97 23.03 79.82 71.55 72.04
Phi-1 89.72 20.69 80.27 88.28 26.17 61.87 82.22 83.63
StarCoder 39.10 73.59 49.26 42.53 23.63 55.23 43.45 43.14
CodeGen2 66.15 54.97 58.30 66.04 23.94 55.24 63.06 63.15
WizCoder 64.41 51.38 57.72 57.91 27.22 36.61 55.29 55.96
CodeLlama 57.79 66.74 59.09 58.60 24.65 53.55 55.54 56.44
Average 64.26 51.01 62.42 64.22 24.77 57.05 61.85 62.39
Table 3. The AUC of feature- and pre-training-based methods in terms of programming languages
Language Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
C / 99.01 99.13
C# / / 98.15
C++ / / 99.58
Go / / 94.19
Java / / 96.11
Python 89.60 / 95.17
Ruby / / 93.25
Rust / / 98.64
When comparing Perplexity against the other two classes of methods, we select the best-performing method as representative [^2] For succinctness, we adopt this measure throughout the paper.
Perplexity does not perform as well as GPTSniffer, which is a pre-training-based method in all eight programming languages. The AUC of Perplexity and GPTSniffer for each language are as follows: C (90.40%–99.13%), C# (90.12%–98.15%), C++ (91.37%–99.58%), Go (72.68%–94.19%), Java (85.54%–96.11%), Python (63.89%–95.17%), Ruby (73.44%–93.25%), and Rust (89.72%–98.64%). The difference can be explained that GPTSniffer is a supervised method while Perplexity is not. Moreover, Perplexity cannot achieve the same performance as feature-based methods in Python and C: Python (63.89%–89.60%), C (90.40%–99.01%).
When comparing the AUC of Perplexity between programming languages, we observe that Perplexity performs best in C/C++ in which NPR and DetectCodeGPT achieve an average AUC of 85%. However, NPR, DetectCodeGPT, and other methods cannot achieve an AUC of over 80% in other languages. Taking C/C++ and Python as examples. C/C++ requires developers to implement many functions, which are more likely to be provided by third-party libraries in Python. In this case, C/C++ has more code stylometry than Python (Caliskan-Islam et al., 2015). However, the design philosophy of Python encourages simplified code stylometry, making it easy for LLMs to generate human-style code.
$\textbf{RQ}_{\textbf{1.1}}$: Perplexity can not perform as well as the feature- and pre-training-based methods from the perspective of programming languages. Among eight languages, Perplexity is only effective in C and C++.
5.2. $\text{RQ}_{\text{1.2}}$: Degree of Difficulty
Table 4 and Table 5 present the AUC of the three groups of methods from the perspective of the degree of difficulty.
Table 4. The AUC of Perplexity in terms of degrees of difficulty
Difficulty Code LLM without perturbation with perturbation
log-p Ent Rank LRank LRR DGPT NPR DCGPT
Easy Incoder 51.75 46.33 40.75 57.76 61.56 46.80 65.03 64.46
Phi-1 69.63 44.47 57.08 69.34 38.90 57.84 75.41 75.09
StarCoder 56.23 52.40 49.72 56.86 50.90 46.27 64.13 62.79
CodeGen2 62.19 46.78 53.82 59.32 49.77 45.71 65.50 64.02
WizCoder 67.84 41.60 55.04 59.52 39.82 47.72 64.72 64.31
CodeLlama 59.84 49.69 53.81 58.15 52.16 45.62 63.60 62.91
Average 61.25 46.88 51.70 60.16 48.85 48.33 66.40 65.60
Middle Incoder 53.30 45.55 41.57 57.04 66.77 42.51 66.02 65.56
Phi-1 76.54 39.33 65.22 71.44 42.02 46.18 76.29 77.00
StarCoder 55.35 56.54 52.41 54.32 49.97 43.12 61.36 60.56
CodeGen2 62.77 49.94 53.28 63.96 53.03 46.77 69.02 68.22
WizCoder 66.38 43.22 56.73 62.90 43.73 43.51 68.18 68.69
CodeLlama 57.67 54.05 52.91 60.93 55.40 41.20 66.08 65.97
Average 62.00 48.11 53.69 61.76 51.82 43.88 67.82 67.67
Hard Incoder 54.53 44.44 37.73 64.05 76.48 41.64 70.27 70.92
Phi-1 75.92 40.94 70.32 72.25 46.92 41.63 74.69 76.07
StarCoder 56.30 57.90 58.33 54.99 53.88 34.26 62.48 62.47
CodeGen2 65.25 50.16 61.54 67.26 57.51 39.70 71.23 71.21
WizCoder 66.28 41.05 63.58 66.32 49.05 36.55 69.21 70.53
CodeLlama 62.97 54.87 65.99 62.32 60.48 35.16 65.82 66.67
Average 63.54 48.23 59.58 64.53 57.39 38.16 68.95 69.64
Table 5. The AUC of feature- and pre-training-based methods in terms of degrees of difficulty
Difficulty Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
Easy 85.37 96.40 90.57
Middle 89.60 98.98 94.32
Hard 93.18 98.25 94.33
Perplexity exhibits lower performance compared with the feature- and pre-training-based methods in terms of all three degrees of difficulty. The differences between Perplexity and pre-training-based methods for each level of difficulty are as follows: easy (75.41%–90.57%), middle (77.00%–94.32%), and hard (76.48%–94.33%). When compared with feature-based methods, Perplexity is less effective than Idialu: easy (75.41%–85.37%), middle (77.00%–89.60%), and hard (76.48%–93.18%), and Bukhari: easy (75.41%–96.40%), middle (77.00%–98.98%), and hard (76.48%–98.25%). The discrepancy can be explained that feature- and pre-training-based methods are based on supervised training, while Perplexity not.
When it comes to the AUC of Perplexity from degrees of difficulty, we find that Perplexity does not perform well in all three degrees of difficulty, but slightly improves from the easy to the hard level. To be specific, log p(x) and Log Rank are the best-performing methods in Perplexity without perturbation, achieving the AUC of 61.25% and 60.10% at the easy level, 62.00%, and 61.76% at the middle level, and 63.54% and 64.53% at the hard level, respectively. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve the AUC of 66.40% and 65.60% at the easy level, 67.82% and 67.67% at the middle level, and 68.95% and 69.64% at the hard level, respectively. Since the solutions to difficult puzzles have complex logic and customized implementations, they have more code stylometry than solutions to easy puzzles.
$\textbf{RQ}_{\textbf{1.2}}$: Perplexity perform not well in all three degrees of difficulty, even though the performance slightly improves from the easy to the hard level. The feature- and pre-training-based methods perform well at all three degrees of difficulty.
5.3. $\text{RQ}_{\text{1.3}}$: Scale of Solution
Table 6 and Table 7 show the AUC of three classes of methods in terms of the scale of solutions. Note that in Table 7, we provide results only in the small and middle scales for Idialu due to no large-scale code solution in Python.
Table 6. The AUC of Perplexity in terms of scales of solution
Scale Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
Small Incoder 52.94 45.41 47.20 52.79 53.91 47.55 58.46 57.24
Phi-1 65.07 42.43 53.11 62.33 42.86 60.14 69.75 68.63
StarCoder 61.72 43.41 49.76 53.42 41.05 54.93 59.29 56.89
CodeGen2 60.65 43.23 50.69 55.45 44.03 47.30 63.75 61.25
WizCoder 66.42 38.99 53.71 63.48 40.26 52.51 69.47 68.01
CodeLlama 61.68 40.38 54.39 60.45 50.67 43.50 67.95 66.35
Average 61.41 42.31 51.48 57.99 45.46 50.99 64.78 63.06
Middle Incoder 57.47 43.11 39.15 61.29 63.56 57.28 72.93 72.04
Phi-1 75.35 38.13 55.44 67.33 35.52 59.11 77.42 77.26
StarCoder 63.84 47.49 54.01 59.42 51.61 56.54 70.22 69.36
CodeGen2 68.56 41.39 51.73 63.92 44.98 56.55 72.55 71.29
WizCoder 72.77 34.01 52.63 63.36 38.64 57.51 73.97 73.64
CodeLlama 63.85 46.76 51.77 62.86 50.66 54.37 70.38 69.56
Average 66.97 41.82 50.79 63.03 47.50 56.89 72.91 72.19
Large Incoder 72.17 29.05 47.88 75.28 77.86 48.97 77.12 78.39
Phi-1 87.81 21.03 73.76 86.00 58.52 50.81 82.22 83.94
StarCoder 67.85 46.41 61.18 67.59 58.67 45.13 68.30 69.46
CodeGen2 78.21 35.07 65.32 77.19 61.70 43.76 75.46 76.62
WizCoder 82.75 21.61 71.10 81.27 57.75 47.70 81.04 82.24
CodeLlama 76.40 42.48 69.43 76.34 64.85 45.54 76.70 77.94
Average 77.53 32.61 64.78 77.28 63.22 46.99 76.81 78.10
Table 7. The AUC of feature- and pre-training-based methods in terms of scales of solution
Scale Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
Small 85.87 91.20 82.74
Middle 92.58 98.26 96.10
Large / 96.67 94.19
Perplexity does not perform as well as the pre-training-based method, i.e., GPTSniffer in all three scales. The AUC of Perplexity and GPTSniffer for each scale is: small (69.75%–82.74%), middle (77.42%–96.10%), and large (87.81%–94.19%). For the featured-based method, Perplexity is less effective than Idialu: small (69.75%–85.87%), middle (77.42%–92.58%), and Bukhari: small (69.75%–91.20%), middle (77.42%–98.26%), and large (87.81%–96.67%). These differences are mainly because feature- and pre-training-based methods are supervised, while Perplexity is a collection of heuristic methods.
When comparing the methods in Perplexity regarding different scales, we observe that those well-performed methods improve from the small to the large scale. For example, log p(x) is the best-performing method in Perplexity without perturbation, with an AUC of 61.41% for small, 66.97% for middle, and 77.53% for large. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve an AUC of 64.87% and 63.06% for small, 72.91% and 72.19% for middle, and 76.81% and 78.10% for large, respectively. These results can be explained that large-scale code snippets have more diverse and detailed implementations than others, which makes HaCode more distinct from LLMgCode.
$\textbf{RQ}_{\textbf{1.3}}$: Perplexity performs less effectively compared with feature- and pre-training-based methods in all the three scales of code solutions. Perplexity only exhibits acceptable performance in the large-scale code solution.
5.4. $\text{RQ}_{\text{2}}$: Detection Speed
Table 8. The detection speed of Perplexity in terms of LLMs (seconds per sample)
Model Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
GPT-3.5 Incoder 0.122 0.092 0.111 0.128 0.235 143.5 5.306 5.306
Phi-1 0.148 0.118 0.137 0.151 0.282 170.7 6.307 6.308
StarCoder 0.273 0.231 0.248 0.267 0.508 152.0 11.12 11.12
CodeGen2 0.328 0.281 0.306 0.319 0.606 148.3 13.27 13.28
WizCoder 0.640 0.561 0.577 0.656 1.265 156.1 30.27 30.27
CodeLlama 0.106 0.076 0.096 0.135 0.240 146.7 5.545 5.545
Average 0.270 0.227 0.246 0.276 0.523 152.9 11.97 11.97
GPT-4o Incoder 0.127 0.099 0.118 0.134 0.248 143.1 5.619 5.620
Phi-1 0.157 0.126 0.146 0.163 0.305 143.5 6.738 6.740
StarCoder 0.292 0.249 0.265 0.289 0.553 148.9 11.97 11.97
CodeGen2 0.360 0.303 0.321 0.350 0.662 151.0 14.29 14.29
WizCoder 0.691 0.604 0.623 0.708 1.379 155.2 33.08 33.08
CodeLlama 0.108 0.080 0.100 0.139 0.248 143.9 5.800 5.800
Average 0.289 0.244 0.262 0.297 0.566 147.6 12.92 12.92
Gemini-1.0 Incoder 0.122 0.094 0.113 0.130 0.238 173.0 5.370 5.370
Phi-1 0.150 0.121 0.140 0.151 0.285 146.1 6.400 6.400
StarCoder 0.278 0.237 0.255 0.267 0.509 169.9 11.14 11.14
CodeGen2 0.336 0.287 0.306 0.328 0.618 146.8 13.37 13.37
WizCoder 0.654 0.575 0.591 0.672 1.288 161.0 30.31 30.31
CodeLlama 0.105 0.078 0.098 0.132 0.237 143.8 5.569 5.570
Average 0.274 0.232 0.251 0.280 0.529 156.8 12.03 12.03
Llama-3.1 Incoder 0.124 0.096 0.114 0.139 0.248 172.0 5.406 5.407
Phi-1 0.153 0.123 0.141 0.155 0.289 165.7 6.479 6.481
StarCoder 0.283 0.241 0.258 0.273 0.522 162.7 11.38 11.38
CodeGen2 0.342 0.292 0.312 0.328 0.622 150.8 13.58 13.58
WizCoder 0.670 0.586 0.603 0.425 0.801 223.7 16.39 16.39
CodeLlama 0.106 0.079 0.099 0.124 0.217 142.6 4.879 4.879
Average 0.280 0.236 0.255 0.241 0.450 169.6 9.685 9.685
Table 9. The detection speed of feature- and pre-training-based methods in terms of LLMs (seconds per sample)
LLM Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
GPT-3.5 0.038 4.352 0.014
GPT-4o 0.042 4.312 0.010
Gemini-1.0 0.040 4.504 0.010
Llama-3.1 0.040 4.794 0.010
Table 8 and Table 9 present the speed of the three classes of methods in detecting code generated by four LLMs. For presentation clarity, we highlight notable results in blue (sec/sample: $0-0.01$;$0.01-0.1$;$0.1-1$)
Perplexity is not as fast as feature- and pre-training-based methods on all four LLMs. The AUC of Perplexity and the pre-training-based method for each LLM are as follows: GPT-3.5 (0.076–0.014), GPT-4o (0.080–0.010), Gemini-1.0 (0.078–0.010), and Llama-3.1 (0.079–0.010). Perplexity calculates the perplexity of each token (a code snippet consists of a set of tokens) using another LLM; therefore, it is very time-consuming. When compared with the feature-based method, Perplexity is slower than Idialu: GPT-3.5 (0.076–0.038), GPT-4o (0.080–0.042), Gemini-1.0 (0.078–0.040), and Llama-3.1 (0.079–0.040). However, Perplexity is faster than Bukhari: GPT-3.5 (0.076–4.352), GPT-4o (0.080–4.312), Gemini-1.0 (0.078–4.504), and Llama-3.1 (0.079–4.794). Since Bukhari extracts n-gram features from abstract syntactic trees, it has a slow detection speed.
When comparing the detection speed of methods in Perplexity, we observe that methods without perturbation are much faster than methods with perturbation. The speed of the worst-performing method without perturbation is less than 2 seconds per sample: GPT-3.5 (1.265), GPT-4o (1.379), Gemini-1.0 (1.288), and Llama-3.1 (0.801). On the contrary, the speed of the best-performing method with perturbation is more than 4 seconds per sample: GPT-3.5 (5.306), GPT-4o (5.619), Gemini-1.0 (5.370), and Llama-3.1 (4.879). Perplexity with perturbations introduces variations into the original sample to generate a list of perturbed samples and calculate attributes such as log-rank of two types of samples, whereas Perplexity without perturbations only calculates attributes of the original sample. Therefore, Perplexity without perturbations performs much faster than Perplexity with perturbations. When comparing Perplexity methods with perturbation, NPR and DetectCodeGPT have a faster detection speed than DetectGPT. The average speed of NPR and DetectCodeGPT is around 10 seconds per sample: GPT-3.5 (11.97–11.97), GPT-4o (12.92–12.92), Gemini-1.0 (12.03–12.03), and Llama-3.1 (9.685–9.685). NPR and DetectCodeGPT achieve very close performance because DetectCodeGPT is based on NPR. However, the average speed of DetectGPT is more than 100 seconds per sample: GPT-3.5 (152.9), GPT-4o (147.6), Gemini-1.0 (156.8), and Llama-3.1 (169.6). The perturbation method of DetectGPT is complex, i.e., masking and completing tokens by another LLM. The perturbation method of DetectCodeGPT is simple, i.e., adding spaces and new lines to the original tokens.
$\textbf{RQ}_{\textbf{2}}$: Perplexity exhibits a lower detection speed than the feature- and pre-training-based methods. The detection speed of Perplexity without perturbation is faster than Perplexity with perturbation.
5.5. $\text{RQ}_{\text{3}}$: Generalization Capability
Table 10 and Table 11 present the AUC of Perplexity and the other two classes of methods in detecting LLMgCode generated by four LLMs. The LLMs listed in the two tables are used to generate LLMgCode. The code LLMs listed in Table 10 are used to calculate perplexity values. For presentation clarity, we highlight notable results in blue (AUC:$90\%-100\%$;$80\%-90\%$;$70\%-80\%$).
Table 10. The AUC of Perplexity in terms of LLMs
LLM Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
GPT-3.5 Incoder 51.89 46.86 39.93 63.01 72.28 49.68 73.10 72.78
Phi-1 75.44 41.93 66.19 75.87 45.27 52.79 81.10 81.64
StarCoder 56.14 54.82 55.86 61.09 57.55 47.53 70.86 70.03
CodeGen2 63.27 49.26 58.56 66.84 56.52 48.27 73.31 72.40
WizCoder 66.34 42.77 60.98 68.54 50.25 72.24 74.26 74.37
CodeLlama 59.41 53.78 60.53 65.79 60.39 41.48 71.52 71.28
Average 62.08 48.24 57.01 66.86 57.04 52.00 74.03 73.75
GPT-4o Incoder 55.29 44.88 44.15 63.16 63.12 54.10 74.89 74.71
Phi-1 72.89 38.02 63.79 72.56 44.10 58.14 80.52 80.88
StarCoder 56.04 52.69 53.07 58.90 53.02 51.63 71.25 70.44
CodeGen2 62.52 48.15 54.12 63.91 51.03 53.03 72.10 71.55
WizCoder 65.54 40.62 57.90 65.91 45.73 66.75 72.95 73.00
CodeLlama 60.70 50.93 56.88 64.68 54.86 48.01 71.13 70.93
Average 62.16 45.88 54.98 64.85 51.98 55.28 73.81 73.59
Gemini-1.0 Incoder 54.53 45.43 42.55 63.70 76.00 51.37 73.74 73.17
Phi-1 74.55 42.25 71.10 74.30 51.08 54.88 81.78 81.66
StarCoder 60.58 53.99 64.18 64.75 66.84 49.24 73.36 72.26
CodeGen2 67.39 48.34 65.84 69.68 65.25 50.68 76.23 74.98
WizCoder 69.21 44.70 64.43 67.99 50.52 58.07 74.36 74.15
CodeLlama 63.76 53.00 68.01 66.36 68.45 43.83 73.49 73.21
Average 65.00 47.95 62.68 67.80 63.02 51.35 75.49 74.90
Llama-3.1 Incoder 57.00 42.88 43.15 65.10 74.82 50.84 73.20 73.14
Phi-1 74.45 41.33 68.64 73.47 46.39 54.96 79.42 79.55
StarCoder 66.20 47.63 67.67 70.29 72.05 51.47 75.56 75.23
CodeGen2 71.42 42.79 66.51 73.05 66.92 49.79 77.42 76.95
WizCoder 74.69 38.59 68.99 76.39 56.50 67.27 79.08 78.74
CodeLlama 69.59 47.23 71.41 73.78 72.37 52.15 77.52 77.69
Average 68.89 43.41 64.39 72.01 64.84 54.41 77.03 76.88
Table 11. The AUC of feature- and pre-training-based methods in terms of LLMs
LLM Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
GPT-3.5 91.43 98.63 92.21
GPT-4o 83.16 98.49 90.30
Gemini-1.0 62.77 92.13 75.21
Llama-3.1 71.76 97.15 83.26
Perplexity has a better generalization capacity than the feature- and pre-training-based methods. The AUC of Perplexity for four LLMs is: GPT-3.5 (81.64%), GPT-4o (80.88%), Gemini-1.0 (81.78%), and Llama-3.1 (79.55%), with a range of 2%. The AUC of pre-training-based methods for each LLM is: GPT-3.5 (92.21%), GPT-4o (90.30%), Gemini-1.0 (75.21%), and Llama-3.1 (83.26%). The max-min difference is over 17%. These differences can be explained that pre-training-based methods are trained on the code generated by GPT-3.5, while Perplexity is LLM-independent. When it comes to feature-based methods, Perplexity outperforms Idialu and Bukhari. The AUC of Idialu for each LLM is as follows: GPT-3.5 (91.43%), GPT-4o (83.16%), Gemini-1.0 (62.77%), and Llama-3.1 (71.76%). The max-min difference is 29%. The AUC of Bukhari for each LLM is as follows: GPT-3.5 (98.63%), GPT-4o (98.49%), Gemini-1.0 (92.13%), and Llama-3.1 (97.15%), with a range of 7%. The max-min range of Bukhari is relatively small. Bukhari only works for C in which HaCode is more different from LLMgCode than other languages, making the detection method easier to detect LLMgCode from HaCode.
When comparing the generalization capability of methods in Perplexity, we observe that methods with perturbation have better generalization than methods without perturbation. Methods with perturbation achieve similar performance on the four LLMs, with the range of average AUC around 3%. For instance, the average AUC of DetectCodeGPT for the four LLMs is: GPT-3.5 (73.75%), GPT-4o (73.59%), Gemini-1.0 (74.90%), and Llama-3.1 (76.88%). In contrast, methods without perturbation do not exhibit such performance, with some methods fluctuating up to 10% on different LLMs. For instance, the average AUC of Rank for the four LLMs is as follows: GPT-3.5 (57.01%), GPT-4o (54.98%), Gemini-1.0 (62.68%), and Llama-3.1 (64.39%). Perplexity without perturbations are based on the absolute perplexity value of the original sample. The absolute perplexity is calculated based on the understanding of the code by the detection model. Therefore, the absolute perplexity varies across code generated by different LLMs, leading to fluctuations in the performance of Perplexity without perturbations. In contrast, Perplexity with perturbations are based on the perplexity discrepancy between the original and perturbed samples, making them more generalized.
$\textbf{RQ}_{\textbf{3}}$: Perplexity exhibits a better generalization capability than feature- and pre-training-based methods. The AUC of supervised methods drops significantly when training and testing sets are generated by different LLMs.
6. Implications
This section discusses the characteristics of LLMgCode and discusses the implications of using Perplexity to detect LLMgCode.
6.1. LLMgCode in Retrospect
The retrospective analysis of LLMgCode primarily focuses on its code length and the resulting impact on the evaluation results.
LLMgCode solutions are inherently more concise compared to HaCode solutions for identical tasks. The LLMgCode solutions and HaCode solutions in our dataset are used to solve the same programming puzzles. However, our dataset has 4,135 HaCode solutions and 8,725 LLMgCode solutions with fewer than 30 LOC. This indicates that LLMgCode solutions are inherently more concise than HaCode solutions. The conciseness of LLMgCode solutions can be explained by three main factors: (1) LLMs are familiar with built-in functions and utility libraries, and prefer using them to simplify implementations, (2) LLMs are trained on large-scale code corpora, and tend to imitate standard and concise coding styles, and (3) LLMs prefer minimal implementations rather than verbose or explanatory styles due to the limitation of tokens.
The conciseness of LLMgCode has a negative impact on the detection accuracy of Perplexity. Perplexity is unsuitable for detecting short code snippets because of its detection mechanism. Since the first few tokens of a code snippet typically have relatively high perplexity due to limited contexts, short solutions are more influenced by this effect than long solutions and therefore tend to have higher perplexity. As a result, Perplexity may misclassify short LLMgCode solutions as HaCode solutions due to their relatively high perplexity. The results of $\text{RQ}_{1.3}$ show that Perplexity does not perform well on small-scale code with an AUC of only 69.75%. As Perplexity is unsuitable for detecting short code snippets, the conciseness of LLMgCode solutions increases the difficulty of detection and reduces the detection accuracy of Perplexity.
To mitigate the impact of code length difference on evaluation results, we evaluate the detection methods from the aspect of code scale in RQ3. Specifically, we evaluate the detection methods on three different code scales, where LLMgCode and HaCode solutions within each scale have comparable lengths. In other evaluation aspects, we do not apply any intervention to limit the code length because the differences in code length arise from the inherent characteristics of code written by LLMs and humans, rather than from our designs.
6.2. For Researchers
Our implications for researchers include both the pros and cons of Perplexity.
Perplexity is unsuitable for high-level programming languages. Perplexity performs well in low-level languages such as C/C++ (91%) but poorly in high-level languages such as Python (64%) and Ruby (73%), which implies that Perplexity has limited capability for high-level languages. High-level languages encourage a concise and clear code style. High-level languages such as Python provide many built-in operations for basic elements, e.g., list and dictionary; low-level languages have to manually manage array, pointer, etc. The programming patterns of high-level languages can be more easily learned by a code generator than those of low-level languages. Moreover, high-level languages have more available third-party libraries than low-level languages. Developers are used to leveraging third-party libraries in their daily development to simplify implementation. Since both humans and LLMs use third-party libraries in generating code, detecting LLMgCode from HaCode using Perplexity in high-level languages is challenging.
Perplexity is unsuitable for short code solutions. Perplexity achieves an AUC of 87.81% for code solutions with more than 50 LOC, but its performance drops to 69.75% for code solutions with less than 20 LOC. This result implies that Perplexity is ineffective for short solutions. Perplexity is based on the probability that an LLM assigns to the next token being the ground-truth token based on previous tokens. Therefore, the perplexity of the first few tokens of code snippets (i.e., a collection of tokens) is unstable due to the lack of contexts. In this case, the perplexity of the first few tokens is relatively high. When the length of code snippets is short, the high perplexity of the first few tokens significantly affects the average perplexity of the entire code snippets. In this case, the perplexities of LLMgCode and HaCode are both high, which makes Perplexity confused.
Perplexity is unsuitable for just-in-time detection. NPR and DetectCodeGPT are the best-performing methods in Perplexity. The detection speeds of NPR and DetectCodeGPT are all around 10 seconds per sample, while the detection speed of the pre-training-based method is less than 0.02 seconds per sample. The difference implies that Perplexity is not suitable for just-in-time detection. This can be explained that Perplexity uses an LLM to calculate the perplexity of each token. The efficiency of LLMs in calculating perplexity for each token is lower than that of pre-trained models in calculating a probability for an entire code snippet. Moreover, NPR and DetectCodeGPT are perturbation-based methods. Such methods calculate the perplexity for one original sample and a collection of perturbed samples. As the number of perturbed samples increases, perturbation-based methods become time-consuming.
Although our investigation indicates that Perplexity has several shortcomings, it exhibits advantages in the following aspects.
Perplexity exhibits the overall good generalization capability. Perplexity with perturbation achieves stable performance in detecting LLMgCode across four LLMs, with AUC variations of around 2%. This implies that Perplexity has a good generalization capability when detecting LLMgCode generated by different LLMs. However, Perplexity without perturbation does not achieve such a performance, e.g., the AUC of LRR varies by more than 10% between LLMs. This highlights the importance of incorporating perturbation when using Perplexity to ensure the generalization capability. The detection mechanism of Perplexity is similar to LLMs for code generation. Therefore, Perplexity remains useful provided that LLMs are based on the auto-regressive generation mechanism (i.e., generating a token based on tokens generated before). The performances of supervised methods are not well when the code snippets for training and testing are generated by different LLMs. The results indicate that supervised methods primarily learn the coding style of LLMs that generate training samples rather than the intrinsic differences between LLMgCode and HaCode. Considering new powerful LLMs are emerging, e.g., DeepSeek, Perplexity is the most suitable method to address this issue.
Perplexity performs well in detecting LLMgCode in C/C++. Perplexity achieves an AUC of 90.40% in C and 92.10% in C++, respectively. This performance is impressive given that Perplexity is a heuristic method rather than a supervised method. These results demonstrate that Perplexity is suitable for detecting LLMgCode in C/C++. C/C++ are low-level programming languages with fewer available third-party libraries than high-level languages such as Python. Therefore, developers implement basic functionality manually. Additionally, C/C++ emphasizes more manual control than high-level languages. Therefore, developers handle details like variable types and dynamic memory. These characteristics make C/C++ code diverse due to personal coding styles and hard for LLMs to imitate. Perplexity performs not as well as the pre-training method in C/C++ LLMgCode of GPT-3.5. However, the pre-training method has poor generalization capability since it is supervised, with an AUC dropping from 92. 21% on GPT-3.5 to 75.21% on Gemini-1.0. In contrast, Perplexity demonstrates the best generalization with the AUC around 80% on all four LLMs, indicating that Perplexity is LLM-independent. Therefore, considering both performance and generalization, we recommend using Perplexity to detect LLMgCode in C/C++.
Perplexity exhibits the good interpretability. Perplexity simulates the code generation mechanisms of LLMs to calculate the perplexity of code snippets. Therefore, the perplexity value is interpretable. A code snippet with a high perplexity indicates that it is disjointed, which might not be LLMgCode. The threshold of perplexity for detection can be adjusted according to the need. Moreover, the calculation of perplexity is performed token by token, allowing it to be precise for the line-level or even token-level detection. This enables Perplexity to perform fine-grained detection. Fine-grained detection does not mean detecting short code. It refers to detecting code snippets in which only a few lines are generated by LLMs. The detection mechanism makes Perplexity applicable to fine-grained detection. For instance, if most of a code snippet is written by a developer and only a few lines are generated by the LLM, Perplexity can calculate the perplexity line by line and detect the lines with low perplexity as LLMgCode. The mechanisms of Perplexity help reviewers understand detection results and provide a foundation for detecting intricate cases where LLMgCode and HaCode are mixed. The mechanisms of supervised methods are complex and hard to interpret. They only output a prediction score or binary result, making developers difficult to understand why the code is detected as LLMgCode and which part is LLMgCode.
In short, there is a trade-off between detection accuracy, speed, and generalization capability. Perplexity is not the most effective in terms of accuracy or efficiency, but it offers the best generalization capability. Unlike supervised methods that learn the characteristics of a specific LLMgCode (e.g., GPT 3.5-generated code), Perplexity seeks to understand the intrinsic differences between LLMgCode and HaCode, making it robust when applied to code generated by unknown LLMs. Although Perplexity is LLM-independent and achieves similar performance across different sources, it has suboptimal accuracy and slower detection speed than feature-based and pre-training-based methods. The best-performing methods of Perplexity are based on perturbation, which introduces additional computational overhead and slows the detection. Therefore, Perplexity is not suitable for just-in-time detection. Separately, Perplexity can be naturally extended to fine-grained detection (e.g., line level) when LLMgCode and HaCode are mixed, whereas other methods provide only a binary classification result (or a prediction score) for the entire code snippet.
Based on the experimental results and the analysis of Perplexity’s detection mechanism, we provide the following directions to improve the detection accuracy and speed of Perplexity.
Effective perplexity calculation. The current detection LLMs used for perplexity calculation have only a few billion parameters, while LLMs used for code generation in practice typically have hundreds of billions of parameters. The large difference in parameters results in significant differences in code understanding and generation, as evidenced by the considerable differences in code solutions generated for the same task. The discrepancy in code understanding and generation between the detection LLM and the generation LLM is a significant reason for the limited effectiveness of Perplexity. To address this problem, first, we recommend using larger LLMs to calculate perplexity, which reduces the differences in perplexity understanding between detection LLMs and generation LLMs, but also increases the computational cost. Second, we recommend distilling the domain knowledge of LLMs with large parameters into LLMs with small parameters. For instance, we can first collect the code solution generated by LLMs with large parameters, e.g., GPT-4o, then utilize fine-tuning or reinforcement learning techniques to train an LLM with small parameters, e.g., CodeLlama-7B, in the domain of code generation, and use the distilled LLMs to calculate the perplexity to detect LLMgCode. By aligning code understanding between detection and generation LLMs, the effectiveness of Perplexity can be improved.
Optimized perturbation method. The perturbation method based on code completion is highly inefficient due to dozens of code masking and completion. Therefore, reducing the number of perturbations is necessary. For instance, instead of randomly masking and completing code lines dozens of times, we can first calculate the perplexity of each line in the original code snippet, and then apply only a few perturbations to the lines with high perplexity. This measure not only improves efficiency several-fold but also amplifies the perplexity discrepancies before and after perturbation, which improves the detection effectiveness. Compared with code completion, adding spaces and empty lines to code snippets is more efficient. However, these perturbation methods are not effective for all types of code, e.g., Python. For indentation-sensitive languages like Python, inserting spaces can cause syntax errors, making the code difficult to understand, i.e., leading to high perplexity for both LLMgCode and HaCode. Therefore, we can further optimize perturbation techniques by considering the characteristics of programming languages. For instance, we can first identify the programming languages of the detected code. If the code is identified as Python, we can apply more targeted perturbation methods, such as replacing variable names. The performance of Perplexity can be improved by implementing optimized perturbation methods.
Improved perplexity utilization. The current utilization of perplexity is insufficient. Representing the perplexity of a code snippet by the average perplexity of each token leads to significant information loss. Two snippets with the same average perplexity could differ in that one has an overall higher perplexity, while the other has a lower overall perplexity but very high perplexity in certain tokens. The perplexity of a code snippet forms a continuous distribution based on the perplexity of each token in sequence. Instead of calculating the average perplexity of each token, we need to better leverage the entire perplexity sequence. For example, we can count the occurrences where the perplexity exceeds a certain threshold, referred to as surpassing occurrences. HaCode tends to be more random and therefore is likely to have more very high perplexity on certain tokens compared to LLMgCode. For another example, we can focus on analyzing changes in the perplexity sequence at the perturbed lines before and after the perturbation. This method amplifies the effect of perturbation, preventing it from being neutralized by the average. The larger the change, the more likely it is LLMgCode, as LLMgCode is more sensitive to perturbations (Shi et al., 2025). Additionally, we can use the perplexity sequence as the initial input for a supervised classifier. Combining the perplexity sequence with supervised learning can improve detection performance compared with Perplexity (heuristic methods).
We propose the following recommendations for the use of Perplexity.
  • Detection method selection:
    • We recommend using DetectCodeGPT or NPR for detection under normal circumstances, as they show the best detection accuracy, acceptable detection speed, and good generalization capability. DetectCodeGPT and NPR perform best among Perplexity, achieving an AUC of around 80% for different LLMs. Moreover, DetectCodeGPT and NPR demonstrate an acceptable detection speed, detecting a sample in just a few seconds.
    • We recommend using Log-Rank for detecting large volumes of samples or for just-in-time detection, as it offers acceptable effectiveness and a fast detection speed. Although Perplexity with perturbation offers better detection performance, it takes several tens of seconds to detect a single sample. Log-Rank performs best among Perplexity without perturbation, with an AUC close to 70% in most cases, while taking only a fraction of a second to detect a sample (50 times faster than DetectCodeGPT/NPR).
  • Application scenario of Perplexity:
    • We recommend applying Perplexity to detect LLMgCode in C/C++ or with a large scale. For the programming language, Perplexity exhibits good performance in C/C++, with an AUC above 90%, while performing suboptimal in other languages like Python, where it only reaches an AUC of 63%. For the code scale, Perplexity exhibits good performance in code solutions with a large scale, achieving an AUC of 87%, while it has limited effectiveness in code solutions with a small or middle scale.
    • We recommend applying Perplexity when the generation models of LLMgCode are unknown, as Perplexity has the best generalization capability. Perplexity achieves an AUC of around 80% regardless of LLMs. However, supervised methods learn the coding styles of LLMs that generate training samples, leading to significant performance fluctuations across different LLMs. Given the continuous emergence of new LLMs and the availability of numerous LLMs for code generation, it is difficult to know whether the training set includes samples generated by the model that generated the code to be detected.
    • We recommend applying Perplexity for fine-grained detection since it can be applied to line-level detection. When a code snippet contains both HaCode and LLMgCode, feature- and pre-training-based methods can only assign a single predictive score to the entire snippet and can not make accurate distinctions. In contrast, Perplexity calculates the perplexity token by token, allowing it to assess the perplexity of each line individually.
6.3. For Tool Development and Usages
When detecting LLMgCode, it is recommended to first detect whether it is LLMgCode and whether it is entirely or partially generated by LLMs, then detect which parts (lines or blocks) are generated by LLMs.
The first step can be formulated as a classification problem. However, the experimental results show that none of the three detection methods (classifiers) are effective. Perplexity only performs well for detecting code in C/C++ (from the perspective of language), and with a large scale (from the perspective of scale). The feature- and pre-training-based methods have high effectiveness but poor generalization capability. Overall, they are all acceptable with an AUC of more than 70%. An ensemble learning approach can fully use each separate method’s advantages while overcoming its disadvantages to some extent. Taking detecting C code as an example, we recommend using DetectCodeGPT (Perplexity), Bukhari (feature-based method), and GPTSniffer (pre-training-based) to develop an ensemble detection tool. By employing an ensemble learning strategy such as MoE (Mixture Of Experts), the detection tool can automatically learn which method(s) to prioritize in different contexts.
The second step is fine-grained detection, i.e., identifying which lines are generated by LLMs. Having understood that the code is generated by LLMs, the second step is to first calculate the perplexity of each token in the code and then calculate the average perplexity for each line. These lines with low perplexity are detected as LLMgCode. We recommend highlighting LLMgCode and HaCode in different colors to remind developers.
The detection tool can be applied in but not limited to following scenarios.
Software supply chain management. LLMgCode has become a new third-party asset in software supply chain development (Bukhari et al., 2023). The widespread use of LLMgCode raises concerns regarding code provenance, accountability, and the potential for introducing bugs or vulnerabilities into software. To address these challenges, the detection tool provides a solution for managing LLMgCode. By detecting the proportion of LLMgCode in third-party libraries, enterprises can assess the reliability of these libraries. By detecting LLMgCode in internal code and centrally managing it as a third-party asset, enterprises can verify compliance with licensing requirements to reduce violations. When there is a problem with the code generated by a certain LLM, it can be quickly located and measures can be taken.
Code review. Since LLMgCode is more prone to include defects (Perry et al., 2024; Sandoval et al., 2023), particular attention is required during the code review process. LLMgCode can lead to unintended risks due to the inherent limitations of LLMs, such as the inclusion of insecure code in the training set (Asare et al., 2023). Therefore, it is crucial for reviewers to be aware of the potential risks associated with LLMgCode and to perform extra scrutiny. The detection tool serves as a valuable aid in this process by automatically detecting and highlighting LLMgCode, ensuring that reviewers are reminded to apply additional checks. These checks may include not only the detection of unique vulnerability patterns targeting LLMgCode, but also plagiarism detection to ensure that LLMgCode is compliant with licensing regulations.
Productivity assessment. Many developers use LLMs to assist with programming, which has led to a shift in how developer productivity is measured (Shi et al., 2025). Traditional metrics, such as LOC, have become less applicable because a significant portion of the submitted code is LLMgCode. LOC-based assessments treat a line of LLMgCode and HaCode as equivalent, resulting in biased evaluations of developers’ actual contributions. To mitigate this challenge, the detection tool helps detect LLMgCode in the submitted code, enabling enterprises to apply more detailed and approximate criteria for assessment. After distinguishing LLMgCode and HaCode, custom evaluations can be performed for each, allowing for a fairer assessment of developer productivity.
Originality assessment. LLMgCode is sometimes not considered original (Nguyen et al., 2024), which raises a concern in contexts that demand independent problem-solving, intellectual property protection, or academic integrity. The detection tool is useful in such situations by automatically determining whether the submitted code is mainly written by humans or generated by LLMs. It can be applied in various scenarios, including but not limited to software competitions, where independent problem-solving skills are valued, and student assignments, where academic honesty and plagiarism prevention are essential. Using this tool, organizations can ensure that submissions meet the necessary standards of originality.
7. Threat to Validity
In this section, we discuss the possible threats to validity and our efforts to mitigate their effects.
Internal Validity: The functions in our dataset are designed to solve specific programming puzzles and are mainly self-contained. This may differ from functions in real-world software repositories, where many customized functions exist and cross-file function calls are common. Although we can extract HaCode from repositories (e.g., GitHub), generating accurate task descriptions from HaCode is difficult without domain knowledge. Without task descriptions, LLMs can not generate precise LLMgCode solutions. Using HaCode and LLMgCode which solve different problems can introduce bias. For example, using HaCode that addresses real-world tasks and LLMgCode that solves programming puzzles may result in LLMgCode being less complex than HaCode, causing bias in the evaluation results. Therefore, we build our dataset based on CodeNet (Puri et al., 2021), which offers a clear task description for each problem and has been widely used in prior work (Xu and Sheng, 2024; Naik et al., 2023; Pan et al., 2024).
Construct Validity: We excluded trivial problems in the dataset as they rarely occur in real-world scenarios. According to  Sheard et al. (2013), there is a strong correlation between code length and problem difficulty. Therefore, we excluded programming problems whose human solutions are all less than 30 LOC to filter out trivial problems. We adopted the thresholds used in prior work–10 (Xu and Sheng, 2024) and 100 (Nguyen et al., 2024)–to determine the initial range for our threshold setting. Within this range, we tested the number of filtered problems under different thresholds at intervals of 10. When the threshold exceeds 50, the number of filtered problems increases rapidly, and this growth becomes steep when the threshold exceeds 80. Excessive filtering introduces risks of excluding non-trivial problems. Therefore, to ensure the generalization of threshold setting, we adopt a safe value of 30. We manually checked the problems filtered by the threshold of 30. All of them are trivial problems or problems without valid submissions.
External Validity: The models used for perplexity calculation range from 1 billion to 7 billion parameters. We examine the statistical correlation between the model size and the evaluation results using Spearman’s rank correlation (Spearman, 1961), which has been widely used in previous work (Niazi, 2015; Hijazi et al., 2022; Lu et al., 2012; Zhou et al., 2009). The resulting p-values are all above 0.05, indicating no statistically significant correlation. This can be explained by the fact that, although the models for perplexity calculation differ in parameter size, they are all much smaller than the code-generating models (such as 175 billion). As a result, their code understanding capabilities are all significantly weaker than those of generation models, resulting in no statistically significant correlation between model size and evaluation results. Small models generally have weaker code understanding capabilities than large models. However, it is impractical for us to locally deploy large generation models (e.g., GPT-3.5) due to resource constraints. Therefore, we use the same model as prior work (Shi et al., 2025) when calculating perplexity.
8. Conclusion
LLMgCode has become increasingly common in daily software development; however, its quality and security remain challenging. Moreover, LLMgCode is sometimes mixed with HaCode. Human-LLM code authorship attribution improves the transparency of software development and increases the trust of third-party software in the supply chain. In this article, we conduct a large-scale investigation of Perplexity, the state-of-the-art method for detecting LLMgCode from HaCode. Toward realistic evaluations, we use 11,664 HaCode snippets and 13,164 LLMgCode snippets to evaluate Perplexity from three perspectives: detection accuracy, detection speed, and generalization capability. The experimental results show that Perplexity is less effective than other methods. Perplexity only performs well in C/C++ from the perspective of language and in a large scale from the perspective of scale. Additionally, Perplexity has a slower detection speed than other methods. However, Perplexity demonstrates strong generalization capability, making it robust in scenarios where the LLM source is unknown. It also has good interpretability and can be naturally applied to fine-grained detection. In summary, there is a trade-off between detection accuracy, speed, and generalization capability: Perplexity achieves similar performance across LLMs but has relatively low accuracy and slow speed. While supervised methods achieve higher accuracy and faster speed, their performance varies when the LLMs for code generation are unknown. In the future, we will continually improve Perplexity to detect LLMgCode from HaCode and evaluate it in industry.
Data Availability
The data and materials of this article are available at https://figshare.com/s/10035be3b64e30e6e779.
Acknowledgments
This work is supported by the Key Research and Development Program of Jiangsu Province (No.BE2021002-2), the Natural Science Foundation of Jiangsu Province (No.BK20241195), the Open Project of Key Laboratory of Industry and Information Technology Ministry for Software Fusion Application and Testing Verification, the National Natural Science Foundation of China (No.62302210, No.62202219), and the Innovation Projects and Overseas Open Projects of State Key Laboratory for Novel Software Technology (Nanjing University) (ZZKT2025A12, ZZKT2025B18, ZZKT2025B20, ZZKT2025B22, KFKT2025A17, KFKT2025A19, KFKT2025A20, KFKT2024A02, KFKT2024A13, KFKT2024A14, KFKT2023A09, KFKT2023A10). Alberto Bacchelli gratefully acknowledges the support of the Swiss National Science Foundation through the SNF Project 200021_197227.